Jefferson Lab HALOG ARCHIVES Logbook for June 09 - All EXCEPT Automatic Entries

  • Main INDEX

    Tuesday

    
    

    Monday

    
    

    Friday

  • 281943 06/26/09 17:17 A. Camsonne Compton Sr90 run collimated run 17683
  • 281942 06/26/09 07:47 Jack (Techs) hallasc18 serial port is OK
  • Wednesday

  • 281941 06/24/09 14:16 A. Camsonne Run number
  • 281940 06/24/09 14:16 A. Camsonne Electron detector noisy modified board
  • 281939 06/24/09 11:35 V. Sulkosky Lent PS757 Logic to PVDIS
  • 281938 06/24/09 09:20 V. Sulkosky BigBite TS Off
  • 281937 06/24/09 09:10 V. Sulkosky Cables to Weldment from HRS Disconnected (late entry)
  • Tuesday

  • 281936 06/23/09 17:13 Jin Huang More NIM modules -> Storage Cabinate
  • 281935 06/23/09 17:06 Jin Huang Target NIM Crate Dissembled
  • Saturday

  • 281934 06/20/09 11:35 A. Camsonne Swipe of Compton chamber OK, will start with source test
  • Friday

  • 281933 06/19/09 18:03 brads BB crates disabled in HVGUI
  • 281932 06/19/09 17:23 Target Peopl no_key
  • 281931 06/19/09 15:47 A. Camsonne Compton electron detector opened for swipe
  • Thursday

  • 281930 06/18/09 12:11 Jack BigBite gas system
  • 281929 06/18/09 11:47 V. Sulkosky BigBite Crates Off
  • 281928 06/18/09 11:23 V. Sulkosky Run 22548 (pedestal run)
  • 281923 06/18/09 11:06 V. Sulkosky Run 3694 (pedestal run)
  • 281920 06/18/09 10:47 Chiranjib/Ya COMPASS SURVEY
  • 281919 06/18/09 09:49 V. Sulkosky Cosmics Runs for LHRS/BB and HAND
  • 281918 06/18/09 09:46 V. Sulkosky BigBite and HAND HV's Off
  • Wednesday

  • 281911 06/17/09 15:16 jpchen scattering chamber swipes results: clean
  • 281910 06/17/09 15:10 X. Zheng HAPPEX run 31554
  • 281909 06/17/09 15:06 Yi Qiang Compass is ready to be surveyed
  • 281904 06/17/09 14:48 X. Zheng PVDIS discriminator thresholds changed (back)
  • 281903 06/17/09 14:22 Yi Qiang Current Settings for Compass survey ( a very tiny little change in one last digit )
  • 281902 06/17/09 14:16 Yi Qiang Current Settings for Compass survey
  • 281901 06/17/09 13:20 X. Zheng stopped HAPPEX run 31553
  • 281900 06/17/09 13:01 R. Michaels re: DAQs
  • 281899 06/17/09 12:35 X. Zheng started HAPPEX run 31553
  • 281894 06/17/09 10:45 Fansler Techs
  • 281889 06/17/09 09:56 V. Sulkosky Cosmic runs for BigBtie and HAND
  • 281885 06/17/09 09:44 jpchen cooling flow off. scattering chamber open/radcon survey
  • 281884 06/17/09 08:59 V. Sulkosky HAND and BigBite HV's Off
  • 281883 06/17/09 08:56 V. Sulkosky HV Screen for BB/HAND/RHRS
  • 281882 06/17/09 08:54 V. Sulkosky Discovered BigBite and HAND HV's are On
  • 281881 06/17/09 05:35 pkpan summary of online replay in owl shift
  • 281880 06/17/09 05:06 A. Camsonne Compton data taking stopped
  • 281879 06/17/09 04:40 auto_snap Happex Charge Asym snapshot
  • 281878 06/17/09 04:40 auto_snap Hall A Tools snapshot
  • 281877 06/17/09 04:40 auto_snap Compton Polarimeter snapshot
  • 281876 06/17/09 04:40 auto_snap BigBox Magnet snapshot
  • 281875 06/17/09 04:40 auto_snap Beam Position and Current snapshot
  • 281874 06/17/09 04:40 auto_snap Accelerator Energies snapshot
  • 281869 06/17/09 04:18 A. Camsonne Reduced HV on photon background rates high
  • 281854 06/17/09 03:21 pkpan VDC trips, and frequent beam trips
  • 281847 06/17/09 03:00 mcnulty BPM/BCM strip charts
  • Tuesday

  • 281824 06/16/09 23:54 dalton Shift Summary
  • 281823 06/16/09 23:19 dalton beam at 100 uA
  • 281811 06/16/09 21:57 R. Michaels Compton activity, part III
  • 281807 06/16/09 21:42 A. Camsonne Analog signal 3 uA at 11 mm
  • 281806 06/16/09 21:17 R. Michaels Compton activity, part II
  • 281801 06/16/09 21:06 A. Camsonne Compton electron detector moved to 11 m
  • 281800 06/16/09 20:40 auto_snap Target Field and Oven Temps snapshot
  • 281798 06/16/09 20:39 A. Camsonne Compton electron detector analog signal at 6 uA
  • 281799 06/16/09 20:39 auto_snap Target Cell Temps snapshot
  • 281797 06/16/09 20:39 auto_snap Radiation and Ozone snapshot
  • 281796 06/16/09 20:39 auto_snap Happex Charge Asym snapshot
  • 281795 06/16/09 20:39 auto_snap Hall A Tools snapshot
  • 281794 06/16/09 20:39 auto_snap Compton Polarimeter snapshot
  • 281793 06/16/09 20:39 auto_snap BigBox Magnet snapshot
  • 281792 06/16/09 20:39 auto_snap Beam Position and Current snapshot
  • 281791 06/16/09 20:39 auto_snap Accelerator Energies snapshot
  • 281790 06/16/09 20:34 A. Camsonne Compton electron detector current at 7 mm 6 uA
  • 281789 06/16/09 20:34 R. Michaels Compton activity, part I
  • 281788 06/16/09 20:24 A. Camsonne Electron detector moved to 7 mm
  • 281783 06/16/09 20:12 R. Michaels Edet pulses and pulseheight at 60 mm
  • 281782 06/16/09 19:52 R. Michaels Typical photon pulses
  • 281777 06/16/09 19:43 R. Michaels FADC pulse heights
  • 281772 06/16/09 18:56 dalton Snack crisis solved
  • 281767 06/16/09 18:24 V. Sulkosky Right Dipole Trip from June 12
  • 281764 06/16/09 18:00 dalton Left HRS
  • 281761 06/16/09 17:35 A. Camsonne ETROC Firmware loaded
  • 281760 06/16/09 17:19 dalton Raster size changed
  • 281747 06/16/09 16:31 A. Camsonne Compton electron detector currents
  • 281742 06/16/09 16:11 A. Camsonne Compton Electron detector currents
  • 281741 06/16/09 16:02 A. Camsonne Background for Compton good
  • 281740 06/16/09 16:01 X. Zheng Runs before 15:20 had HWP OUT
  • 281739 06/16/09 16:00 X. Zheng Day shift summary
  • 281735 06/16/09 15:36 X. Zheng Beam HWP inserted
  • 281728 06/16/09 15:03 V. Sulkosky Beam for the Past Hour
  • 281723 06/16/09 14:53 Higinbotham BigBite Cards
  • 281722 06/16/09 14:51 X. Zheng LHRS set to "leftHRS" now
  • 281710 06/16/09 13:48 A. Camsonne Compton clean up
  • 281709 06/16/09 13:46 A. Camsonne Beam off for a few minutes for RF work to high current
  • 281692 06/16/09 13:08 V. Sulkosky Re: LHRS DAQ scalers not cleared
  • 281687 06/16/09 12:39 auto_snap Target Field and Oven Temps snapshot
  • 281686 06/16/09 12:39 auto_snap Target Cell Temps snapshot
  • 281685 06/16/09 12:39 auto_snap Radiation and Ozone snapshot
  • 281684 06/16/09 12:39 auto_snap Happex Charge Asym snapshot
  • 281683 06/16/09 12:39 auto_snap Hall A Tools snapshot
  • 281682 06/16/09 12:39 auto_snap Compton Polarimeter snapshot
  • 281681 06/16/09 12:39 auto_snap BigBox Magnet snapshot
  • 281680 06/16/09 12:39 auto_snap Beam Position and Current snapshot
  • 281679 06/16/09 12:38 auto_snap Accelerator Energies snapshot
  • 281678 06/16/09 12:29 A. Camsonne Analog signal plan A in garage position
  • 281677 06/16/09 12:19 A. Camsonne Compton electron detector currents when in garage position and ion chamber check
  • 281676 06/16/09 12:13 A. Camsonne First 6 GeV CW beam in Compton
  • 281673 06/16/09 11:57 X. Zheng target helicity info "fixed"
  • 281672 06/16/09 11:56 Yi Qiang Both 3He spin signals are set to logic 1
  • 281671 06/16/09 11:55 X. Zheng LHRS DAQ scalers not cleared
  • 281664 06/16/09 11:35 A. Camsonne Hall A BCM showing complete 0
  • 281663 06/16/09 11:22 X. Zheng update (BCM chart)
  • 281659 06/16/09 11:03 X. Zheng better spot++
  • 281654 06/16/09 10:57 X. Zheng spot++ MCC unit 6x6.4
  • 281653 06/16/09 10:45 X. Zheng spot++ with MCC 5x5
  • 281638 06/16/09 10:31 X. Zheng target moved to carbon
  • 281637 06/16/09 09:26 R. Michaels Test2 config for Compton
  • 281635 06/16/09 07:11 R. Subedi Status Update
  • 281634 06/16/09 04:38 auto_snap Target Field and Oven Temps snapshot
  • 281633 06/16/09 04:38 auto_snap Target Cell Temps snapshot
  • 281632 06/16/09 04:38 auto_snap Radiation and Ozone snapshot
  • 281631 06/16/09 04:38 auto_snap Happex Charge Asym snapshot
  • 281630 06/16/09 04:38 auto_snap Hall A Tools snapshot
  • 281629 06/16/09 04:38 auto_snap Compton Polarimeter snapshot
  • 281628 06/16/09 04:38 auto_snap BigBox Magnet snapshot
  • 281627 06/16/09 04:38 auto_snap Beam Position and Current snapshot
  • 281626 06/16/09 04:38 auto_snap Accelerator Energies snapshot
  • 281625 06/16/09 04:13 R. Subedi Beam Status Update
  • 281624 06/16/09 00:00 R. Subedi Beam Update
  • Monday

  • 281623 06/15/09 22:42 R. Subedi Called off controlled access
  • 281622 06/15/09 21:06 A. Camsonne Compton electron detector work
  • 281621 06/15/09 21:00 A. Camsonne Compton electron detector moved to garage position
  • 281620 06/15/09 20:38 auto_snap Target Field and Oven Temps snapshot
  • 281619 06/15/09 20:38 auto_snap Target Cell Temps snapshot
  • 281618 06/15/09 20:38 auto_snap Radiation and Ozone snapshot
  • 281617 06/15/09 20:38 auto_snap Happex Charge Asym snapshot
  • 281616 06/15/09 20:38 auto_snap Hall A Tools snapshot
  • 281615 06/15/09 20:38 auto_snap Compton Polarimeter snapshot
  • 281614 06/15/09 20:38 auto_snap BigBox Magnet snapshot
  • 281613 06/15/09 20:37 auto_snap Beam Position and Current snapshot
  • 281612 06/15/09 20:37 auto_snap Accelerator Energies snapshot
  • 281611 06/15/09 20:20 V. Sulkosky Re: MCC update
  • 281610 06/15/09 20:00 Higinbotham Wien Angle
  • 281609 06/15/09 18:21 Jack (Techs) Bogie controls
  • 281608 06/15/09 18:19 X. Zheng MCC update
  • 281607 06/15/09 18:15 Jack hatsv3 on Right Detector Stack was replaced
  • 281606 06/15/09 18:00 R. Michaels re: ROC9
  • 281600 06/15/09 17:19 Ramesh/Xiaoc pvdis access activities (late entry)
  • 281599 06/15/09 17:11 deng target moved to empty
  • 281598 06/15/09 17:02 X. Zheng LHRS DAQ set to "LeftOnly"
  • 281595 06/15/09 16:50 X. Zheng update
  • 281594 06/15/09 16:23 A. Camsonne Compton electron analog signal
  • 281593 06/15/09 16:22 X. Zheng ROC9 problem
  • 281583 06/15/09 15:58 deng happex run 31538 ended, test run
  • 281579 06/15/09 15:37 deng happex run 31537 ended, cosmic run
  • 281578 06/15/09 15:20 A. Camsonne MPS signal in ETROC
  • 281576 06/15/09 15:03 V. Sulkosky L-HRS -> 14 degs and R-HRS -> 18 degs
  • 281573 06/15/09 13:34 gjin setting MWDC thresholds to 0
  • 281572 06/15/09 13:04 Yi Qiang Oven Air Flow and Heaters OFF
  • 281571 06/15/09 13:01 X. Zheng HRS momentum set
  • 281570 06/15/09 12:37 auto_snap Target Field and Oven Temps snapshot
  • 281569 06/15/09 12:37 auto_snap Target Cell Temps snapshot
  • 281568 06/15/09 12:37 auto_snap Radiation and Ozone snapshot
  • 281567 06/15/09 12:37 auto_snap Happex Charge Asym snapshot
  • 281566 06/15/09 12:37 auto_snap Hall A Tools snapshot
  • 281565 06/15/09 12:37 auto_snap Compton Polarimeter snapshot
  • 281564 06/15/09 12:37 auto_snap BigBox Magnet snapshot
  • 281563 06/15/09 12:37 auto_snap Beam Position and Current snapshot
  • 281562 06/15/09 12:37 auto_snap Accelerator Energies snapshot
  • 281561 06/15/09 12:23 Yi Qiang Target Photos
  • 281560 06/15/09 12:13 Higinbotham Right Arm Angle (a.k.a. The Poor Man's Survey)
  • 281559 06/15/09 11:59 Higinbotham BigBite Blast Shield
  • 281558 06/15/09 11:35 V. Sulkosky Experiment Summary for June 14 owl - June 15 owl
  • 281557 06/15/09 11:27 Yi^2 3He cell/Ref cell/Target Ladder removed, Parity target installed
  • 281556 06/15/09 09:53 X. Zheng Changing left and right HRS momentum
  • 281555 06/15/09 09:31 V. Sulkosky Planned Activities During Access
  • 281554 06/15/09 09:27 X. Zheng Hall is in controlled access
  • 281553 06/15/09 09:24 V. Sulkosky Hall Being Surveryed for Controlled Access
  • 281552 06/15/09 09:05 Yi Qiang Lasers are LOCKED
  • 281551 06/15/09 08:29 pkpan owl shift summary (improved version)
  • 281549 06/15/09 08:09 miham Oven Air Flow SP4 set to 800
  • 281548 06/15/09 08:03 pkpan owl shift summary
  • 281547 06/15/09 07:55 miham Target Lasers are OFF
  • 281546 06/15/09 07:34 V. Sulkosky Pedestal Runs 3635/22490
  • 281541 06/15/09 07:28 miham Reference cell vented (Formated)
  • 281540 06/15/09 07:27 miham Reference cell vented
  • 281539 06/15/09 07:22 miham Doug' dog is alive again and feeling well
  • 281538 06/15/09 07:11 miham BigBite and ND HV turned OFF
  • 281537 06/15/09 07:05 Kai/Miha Beam has been taken away (for good)
  • 281536 06/15/09 07:03 luhj NMR55.825%
  • 281529 06/15/09 06:51 adaq spot++ for Run 3633
  • 281524 06/15/09 06:13 A. Camsonne Compton stand alone electron run 17510
  • 281519 06/15/09 05:31 pkpan For log 281495, the old Campton Run was 17505, not 17506, that is ending failed
  • 281510 06/15/09 04:37 auto_snap Target Field and Oven Temps snapshot
  • 281509 06/15/09 04:37 auto_snap Target Cell Temps snapshot
  • 281508 06/15/09 04:37 auto_snap Radiation and Ozone snapshot
  • 281507 06/15/09 04:37 auto_snap Happex Charge Asym snapshot
  • 281506 06/15/09 04:36 auto_snap Hall A Tools snapshot
  • 281505 06/15/09 04:36 auto_snap Compton Polarimeter snapshot
  • 281504 06/15/09 04:36 auto_snap BigBox Magnet snapshot
  • 281503 06/15/09 04:36 auto_snap Beam Position and Current snapshot
  • 281502 06/15/09 04:36 auto_snap Accelerator Energies snapshot
  • 281497 06/15/09 04:19 pkpan For log 281495, the old Campton Run was 17505, not 17506, that is ending failed
  • 281496 06/15/09 04:10 pkpan a new Happex Run 31536 before HRS Run #3628/ #22483
  • 281495 06/15/09 04:08 pkpan old Compton Run 17506 ending failed, reset and start a new Compton Run 17506
  • 281482 06/15/09 02:31 pkpan NMR 56.7%
  • 281473 06/15/09 01:55 ejensen Run #22478 ND replay
  • 281470 06/15/09 01:33 ejensen Wire chamber trips, ended runs
  • 281458 06/15/09 00:01 LeRose Shift Summary
  • Sunday

  • 281445 06/14/09 22:35 Yawei NMR 57.1%, laser absorption spectrum
  • 281430 06/14/09 21:09 LeRose Beam Back
  • 281429 06/14/09 20:52 LeRose Hall A Tools
  • 281428 06/14/09 20:36 auto_snap Target Field and Oven Temps snapshot
  • 281427 06/14/09 20:36 auto_snap Target Cell Temps snapshot
  • 281426 06/14/09 20:36 auto_snap Radiation and Ozone snapshot
  • 281425 06/14/09 20:36 auto_snap Happex Charge Asym snapshot
  • 281424 06/14/09 20:36 auto_snap Hall A Tools snapshot
  • 281423 06/14/09 20:36 auto_snap Compton Polarimeter snapshot
  • 281422 06/14/09 20:36 auto_snap BigBox Magnet snapshot
  • 281421 06/14/09 20:36 auto_snap Beam Position and Current snapshot
  • 281420 06/14/09 20:36 auto_snap Accelerator Energies snapshot
  • 281419 06/14/09 20:13 LeRose No Beam
  • 281415 06/14/09 20:01 Yawei half wave plate in, restart happex run#31534
  • 281409 06/14/09 19:43 Yawei New happex run started, run#31533
  • 281402 06/14/09 18:52 chench Left DAQ crashed, grun=3609(L) and grun=22464(R) are junk
  • 281398 06/14/09 18:46 Yawei NMR Measurement 55.5% Transverse -, laser absorption spectrum
  • 281393 06/14/09 18:27 V. Sulkosky R-HRS DAQ Crashed (22462)
  • 281390 06/14/09 18:18 A. Camsonne Trigger mask for ETROC plan C
  • 281389 06/14/09 18:16 Yawei Spot++ check, reference cell
  • 281385 06/14/09 18:12 V. Sulkosky A few HV Channels Tripped
  • 281384 06/14/09 18:12 Yawei Target moved to Reference Cell Position
  • 281383 06/14/09 18:11 Yawei spot++ check
  • 281378 06/14/09 18:05 A. Camsonne Compton Electron detector work
  • 281377 06/14/09 17:50 Yawei Target moved to Optics
  • 281376 06/14/09 17:06 Yawei Fill D2 into reference cell, 135psig
  • 281375 06/14/09 17:01 Yi Qiang D2 bottle connected to the reference cell
  • 281372 06/14/09 16:25 Yawei NMR measurement: 56.3%, Laser Absorption Spectrum
  • 281367 06/14/09 16:07 Lamiaa Empty Epics variables histograms for run# 3603
  • 281365 06/14/09 16:03 Lindgren Shift Summary
  • 281364 06/14/09 16:01 V. Sulkosky Experiment Summary for June 13 owl - June 13 swing
  • 281363 06/14/09 15:58 puckett One more absorption spectrum
  • 281358 06/14/09 15:22 Lindgren HAPPEX
  • 281349 06/14/09 14:33 V. Sulkosky New Golden Root File (#22455) for R-HRS Physics Replay
  • 281340 06/14/09 13:41 lindgren spot++
  • 281339 06/14/09 13:27 Lindgren Production running
  • 281330 06/14/09 12:39 miham New Golden Root Files (#3596/#22451)
  • 281329 06/14/09 12:38 puckett NMR frequency sweep 56.7% at 12:26, first HALOG attempt failed
  • 281328 06/14/09 12:36 auto_snap Target Field and Oven Temps snapshot
  • 281327 06/14/09 12:36 auto_snap Target Cell Temps snapshot
  • 281326 06/14/09 12:36 auto_snap Radiation and Ozone snapshot
  • 281325 06/14/09 12:36 auto_snap Happex Charge Asym snapshot
  • 281324 06/14/09 12:35 auto_snap Hall A Tools snapshot
  • 281323 06/14/09 12:35 auto_snap Compton Polarimeter snapshot
  • 281322 06/14/09 12:35 auto_snap BigBox Magnet snapshot
  • 281321 06/14/09 12:35 auto_snap Beam Position and Current snapshot
  • 281320 06/14/09 12:35 auto_snap Accelerator Energies snapshot
  • 281319 06/14/09 12:33 puckett Absorption spectrum
  • 281315 06/14/09 12:27 A. Camsonne Timing Jitter Compton modified electron board
  • 281313 06/14/09 12:25 A. Camsonne Timing Jitter Compton modified electron board
  • 281311 06/14/09 12:05 lindgren scalers
  • 281306 06/14/09 11:52 lindgren HV
  • 281305 06/14/09 11:51 Lindgren heliuml evels strip
  • 281304 06/14/09 11:48 Lindgren Beam Position strip
  • 281303 06/14/09 11:45 Lindgren tools and strip beam current
  • 281302 06/14/09 11:29 Lindgren HAPPEX
  • 281297 06/14/09 11:16 A. Camsonne ETROC firmware changed and stand alone run 17469
  • 281284 06/14/09 09:42 Lindgren HV Crash/Beam excursions
  • 281279 06/14/09 09:01 Lindgren day shift starting/BB HV crash
  • 281274 06/14/09 08:49 miham Calibration runs #3584-#3587: Short summary
  • 281269 06/14/09 08:31 Lamiaa Spot++ check.
  • 281268 06/14/09 08:30 miham Reference cell changes - Reference cell is now filled with N2
  • 281265 06/14/09 08:14 puckett NMR frequency sweep = 55.52%
  • 281261 06/14/09 08:00 pkpan owl shift summary
  • 281258 06/14/09 07:38 pkpan a new Happex Run 31530.
  • 281255 06/14/09 07:36 pkpan a new Compton Run 17467
  • 281254 06/14/09 07:23 Yi Qiang Target spin rotated to Transverse - (270 degrees) / NMR:
  • 281243 06/14/09 06:16 miham Target moved to Ref. Cell position
  • 281236 06/14/09 05:40 V. Sulkosky Re: Raster checkout summary
  • 281235 06/14/09 05:36 Luhj Replay Run3575 and 3574 failed
  • 281234 06/14/09 05:29 pkpan old Compton Run # 17465 ending failed, starting a new Compton Run # 17466
  • 281231 06/14/09 05:21 luhj NMR63.145
  • 281228 06/14/09 04:59 adaq beam back after 10 mins shot shut down, and wire chamber seems working now
  • 281225 06/14/09 04:48 pkpan frequent trips during this time, and 15 minutes beam down already
  • 281224 06/14/09 04:45 pkpan wire chamber trips again
  • 281223 06/14/09 04:35 auto_snap Target Field and Oven Temps snapshot
  • 281222 06/14/09 04:35 auto_snap Target Cell Temps snapshot
  • 281221 06/14/09 04:35 auto_snap Radiation and Ozone snapshot
  • 281220 06/14/09 04:35 auto_snap Happex Charge Asym snapshot
  • 281219 06/14/09 04:35 auto_snap Hall A Tools snapshot
  • 281218 06/14/09 04:35 auto_snap Compton Polarimeter snapshot
  • 281217 06/14/09 04:35 auto_snap BigBox Magnet snapshot
  • 281216 06/14/09 04:35 auto_snap Beam Position and Current snapshot
  • 281215 06/14/09 04:35 auto_snap Accelerator Energies snapshot
  • 281208 06/14/09 04:01 pkpan Beam current is nearly 15uA, ask MCC to decrease it to 14uA
  • 281207 06/14/09 03:56 miham Raster checkout summary
  • 281206 06/14/09 03:53 miham Spot++ check (#3579 - Prod.)
  • 281205 06/14/09 03:32 miham Scaler snap shots (#3579/#22434)
  • 281204 06/14/09 03:30 miham Current increased to 14uA
  • 281199 06/14/09 03:25 miham Spot++ check, Raster 5x4MCC (#3578)
  • 281196 06/14/09 03:18 pkpan a new compton run 17465
  • 281195 06/14/09 03:17 miham Spot++ check , Raster 3x2MCC (#3577)
  • 281190 06/14/09 03:10 miham Spot++ check, Raster 7x5MCC (#3576)
  • 281183 06/14/09 02:45 pkpan spot++ for run #3574 and #3575. Raster size 10*8. Hitting in corner!!!
  • 281182 06/14/09 02:44 miham Scaler snap shots (#3575)
  • 281177 06/14/09 02:32 luhj L-HRS deadtime 64%
  • 281174 06/14/09 02:24 luhj run22428,t3=56Hz
  • 281171 06/14/09 02:22 luhj #3573 , trigger
  • 281170 06/14/09 02:16 pkpan new Happex Run 31528 (stop) and 31529
  • 281164 06/14/09 01:30 luhj NMR63.1632
  • 281163 06/14/09 01:27 Luhj Target Moved to He3
  • 281161 06/14/09 01:23 luhj #22426 Trigger
  • 281159 06/14/09 01:22 luhj 3571#trigger
  • 281158 06/14/09 01:19 pkpan spot++ for Run #3571, size is OK
  • 281151 06/14/09 00:44 luhj Move target to Optics
  • 281149 06/14/09 00:23 gjin Swing shift summary
  • 281148 06/14/09 00:12 gjin swing shift summary
  • Saturday

  • 281147 06/13/09 23:16 A. Camsonne Test ETROC firmware trigger trigger_ctrl_00
  • 281146 06/13/09 22:19 puckett ended runs 3569/22424 due to beam down
  • 281143 06/13/09 22:08 puckett Further HV increase on ND veto left channel 21 improves spectra
  • 281142 06/13/09 21:56 gjin ND HV changed to -1300 for L1.9
  • 281137 06/13/09 21:44 Higinbotham Monday's Target
  • 281136 06/13/09 21:29 gjin ND L1.9 HV changed back to -1250V
  • 281131 06/13/09 21:24 miham Bad jumper fibre pictures
  • 281130 06/13/09 21:12 A. Puckett After change of HV, TDC spectrum of ND veto left channel 21 gets worse...
  • 281129 06/13/09 21:01 puckett Runs 22421/3566 ended on EB1 crash
  • 281128 06/13/09 21:00 Yi Qiang ND V-22L HV adjusted to -1150V
  • 281127 06/13/09 20:59 Yi Qiang New Happex run 31527 started before run 3567/22422
  • 281122 06/13/09 20:35 auto_snap Target Field and Oven Temps snapshot
  • 281121 06/13/09 20:35 auto_snap Target Cell Temps snapshot
  • 281120 06/13/09 20:35 auto_snap Radiation and Ozone snapshot
  • 281119 06/13/09 20:34 auto_snap Happex Charge Asym snapshot
  • 281118 06/13/09 20:34 auto_snap Hall A Tools snapshot
  • 281117 06/13/09 20:34 auto_snap Compton Polarimeter snapshot
  • 281116 06/13/09 20:34 auto_snap BigBox Magnet snapshot
  • 281115 06/13/09 20:34 auto_snap Beam Position and Current snapshot
  • 281114 06/13/09 20:34 auto_snap Accelerator Energies snapshot
  • 281109 06/13/09 20:06 gjin halla tools
  • 281108 06/13/09 19:57 gjin spot++ 4x3
  • 281103 06/13/09 19:33 V. Sulkosky Experiment Summary for June 10 owl - June 10 swing
  • 281098 06/13/09 19:20 puckett the "problem" in question also appears in runs before the DAQ crash
  • 281097 06/13/09 18:58 gjin NMR 62.25%
  • 281092 06/13/09 18:55 V. Sulkosky Experiment Summary for June 11 owl - June 12 swing
  • 281091 06/13/09 18:55 A. Camsonne Stand alone run test run 17452
  • 281090 06/13/09 18:53 V. Sulkosky Re: Veto plane left raw TDC 21
  • 281089 06/13/09 18:44 puckett Veto plane left raw TDC 21
  • 281088 06/13/09 18:31 puckett Runs 3561/22416 JUNK-->DAQ crash
  • 281079 06/13/09 18:21 A. Camsonne Run Compton standalone 17449 and more test modified clock
  • 281078 06/13/09 17:58 A. Camsonne Compton stand alone run and new firmware
  • 281069 06/13/09 17:05 gjin changed beam energy settings in replay histogram display
  • 281060 06/13/09 16:14 Yi Qiang Run 3556/22411 have half-wave plate OUT
  • 281058 06/13/09 16:00 Gao Day shift summary
  • 281057 06/13/09 16:00 Yi Qiang Half-wave plate OUT
  • 281046 06/13/09 14:58 pkpan NMR 62.6%
  • 281031 06/13/09 12:49 H, Gao Right HRS at 17 degree
  • 281028 06/13/09 12:34 auto_snap Radiation and Ozone snapshot
  • 281027 06/13/09 12:34 auto_snap Happex Charge Asym snapshot
  • 281026 06/13/09 12:34 auto_snap Hall A Tools snapshot
  • 281025 06/13/09 12:34 auto_snap Compton Polarimeter snapshot
  • 281024 06/13/09 12:34 auto_snap BigBox Magnet snapshot
  • 281023 06/13/09 12:34 auto_snap Beam Position and Current snapshot
  • 281022 06/13/09 12:34 auto_snap Accelerator Energies snapshot
  • 281014 06/13/09 11:51 Yi Qiang Right HRS Dipole Field and Current History after Pass Change
  • 281013 06/13/09 11:36 Yi Qiang Right HRS Dipole Field and Current History
  • 281012 06/13/09 11:21 A. Camsonne Compton run stand alone mode
  • 281007 06/13/09 10:59 pkpan spot++: raster size 3.5*3.5
  • 281006 06/13/09 10:49 pkpan NMR 62.2%
  • 280985 06/13/09 08:04 Ibrahim Owl Shift Summary (June 13, 2009)
  • 280984 06/13/09 08:03 miham Target Strip Charts at the end of the OWL shift
  • 280978 06/13/09 07:36 Ibrahim Hall A Tools (Right Dipole Status)
  • 280974 06/13/09 07:22 miham No Significant change in rates after the raster change
  • 280973 06/13/09 07:16 miham New Golden Run (#3540) for BB and HRS-L
  • 280972 06/13/09 07:02 miham Scaler snap shots after the raster change (#3540)
  • 280971 06/13/09 06:57 miham Spot++ check after the raster change (#3540)
  • 280968 06/13/09 06:33 A. Camsonne Power cycled ETROC
  • 280967 06/13/09 06:23 miham Scaler snap shots (#3539)
  • 280966 06/13/09 06:16 miham NMR Frequency Sweep ( New Polarization is 63.3%) and absorption spectrum
  • 280963 06/13/09 06:09 miham Opposite change in rates after the Raster and Current Change
  • 280960 06/13/09 05:13 Ibrahim Started the Compton DAQ at 3:53 am. New Run 17414
  • 280959 06/13/09 05:11 Ibrahim Restarted the HAPPEX DAQ. New Run 31524
  • 280956 06/13/09 04:44 miham Problematic Laser absorption spectum
  • 280955 06/13/09 04:39 miham Spot++ check after raster change (#3536)
  • 280954 06/13/09 04:34 miham ND, BB HV trip
  • 280952 06/13/09 04:34 auto_snap Target Field and Oven Temps snapshot
  • 280951 06/13/09 04:34 auto_snap Target Cell Temps snapshot
  • 280950 06/13/09 04:34 auto_snap Radiation and Ozone snapshot
  • 280949 06/13/09 04:34 auto_snap Happex Charge Asym snapshot
  • 280948 06/13/09 04:33 auto_snap Hall A Tools snapshot
  • 280947 06/13/09 04:33 auto_snap Compton Polarimeter snapshot
  • 280946 06/13/09 04:33 auto_snap BigBox Magnet snapshot
  • 280945 06/13/09 04:33 auto_snap Beam Position and Current snapshot
  • 280944 06/13/09 04:33 auto_snap Accelerator Energies snapshot
  • 280943 06/13/09 04:28 miham Spot++ check before raster change (#3535)
  • 280942 06/13/09 04:27 Lamiaa Restarted new Happex/Compton runs# 31524/17414
  • 280940 06/13/09 04:14 Ibrahim Spot++ for Run 2533
  • 280939 06/13/09 04:13 miham Laser No3. not working properly
  • 280937 06/13/09 03:56 Ibrahim Rates for He3, 4x4 mm at 2 and 10 uA
  • 280934 06/13/09 03:46 miham Target moved back to pol. 3He position
  • 280933 06/13/09 03:46 V. Sulkosky R-HRS Q3 Cycled and Set
  • 280931 06/13/09 03:44 miham Target lasers turned OFF/ON
  • 280929 06/13/09 03:36 folts tech/Quench reset
  • 280928 06/13/09 03:26 miham Target moved to Optics position
  • 280926 06/13/09 02:37 Lamiaa Restarted new Happex run# 31523
  • 280925 06/13/09 02:32 Ibrahim Restarted the HAPPEX DAQ. New Run 31523
  • 280922 06/13/09 01:59 V. Sulkosky Right Dipole and Q3 Have Faults
  • 280921 06/13/09 01:50 miham Target screens before production
  • 280920 06/13/09 01:49 V. Sulkosky Magnet Levels and Recovery
  • 280918 06/13/09 01:47 miham NMR Frequency Sweep ( New Polarization is 62.63%) and Absorption spectrum
  • 280917 06/13/09 01:45 Ibrahim Spot++ for Run 3529
  • 280915 06/13/09 01:43 miham Multiple trips in ND, BB and HRS-R VDC
  • 280914 06/13/09 01:41 Ibrahim Rates for He3 4x4 mm at 2 and 10 uA
  • 280912 06/13/09 01:30 miham Target moved back to Pol. 3He position
  • 280910 06/13/09 01:26 miham Lasers turned OFF and ON
  • 280908 06/13/09 01:22 V. Sulkosky Left HRS Back Up
  • 280907 06/13/09 01:20 miham Target moved to Optics target
  • 280905 06/13/09 00:05 Gao Swing shift summary
  • Friday

  • 280904 06/12/09 23:51 Jin Huang LHRS & RHRS Coolant Low
  • 280903 06/12/09 23:50 Jin Huang RHRS Dipole Alarm
  • 280892 06/12/09 22:37 V. Sulkosky Spectrometer Motion Problems (Late Entry)
  • 280887 06/12/09 22:13 A. Camsonne Compton vertical scan
  • 280882 06/12/09 21:43 Jin Huang NMR 64.5; before production
  • 280881 06/12/09 21:43 A. Camsonne Compton electron detector checks
  • 280880 06/12/09 21:30 Jin Huang spot ++ on He3, 4x4mm
  • 280873 06/12/09 21:21 Jin Huang spot++; Carbon. 7x5mcc
  • 280864 06/12/09 21:03 A. Camsonne Compton electron detector access and moved in Compton position
  • 280863 06/12/09 20:38 Jin Huang NMR 63.3%; upon control access
  • 280862 06/12/09 20:33 auto_snap Target Field and Oven Temps snapshot
  • 280861 06/12/09 20:33 auto_snap Target Cell Temps snapshot
  • 280860 06/12/09 20:33 auto_snap Radiation and Ozone snapshot
  • 280859 06/12/09 20:33 auto_snap Happex Charge Asym snapshot
  • 280858 06/12/09 20:33 auto_snap Hall A Tools snapshot
  • 280857 06/12/09 20:33 auto_snap Compton Polarimeter snapshot
  • 280856 06/12/09 20:33 auto_snap BigBox Magnet snapshot
  • 280855 06/12/09 20:33 auto_snap Beam Position and Current snapshot
  • 280854 06/12/09 20:33 auto_snap Accelerator Energies snapshot
  • 280853 06/12/09 20:28 A. Camsonne Requested control access to check Compton Electron detector
  • 280847 06/12/09 19:43 xqian Neutron arm TDC saturation?
  • 280838 06/12/09 18:49 Jin Huang Late Entry: Spot++ @ 6x6 and 4x4
  • 280829 06/12/09 18:03 Jin Huang NMR 64.4%
  • 280826 06/12/09 18:00 Jin Huang Spot++ 2x2mm @ he3
  • 280821 06/12/09 17:50 Jin Huang Target -> He3 Polarized
  • 280820 06/12/09 17:48 Jin Huang Spot++; 7x5 MCC
  • 280817 06/12/09 17:46 JinGe/Miha BB & ND HVs turned on
  • 280814 06/12/09 17:39 Jin Huang Target -> Optics
  • 280813 06/12/09 17:35 Jin Huang Harp Scan Result
  • 280811 06/12/09 17:14 H. Gao day shift summary
  • 280810 06/12/09 16:56 Jin Huang Target-> Empty; Laser back on
  • 280809 06/12/09 16:53 Jin Huang Target -> Optics; Laser off
  • 280808 06/12/09 16:52 Yi Qiang Those two peaks are from 4He in the target chamber
  • 280807 06/12/09 16:00 A. Camsonne Compton electron detector noisy
  • 280806 06/12/09 15:49 miham Quick Carbon Optics run Analysis
  • 280805 06/12/09 14:52 miham BB and ND HV turned OFF
  • 280804 06/12/09 14:21 pkpan target moved from He3 to Empty position
  • 280803 06/12/09 12:46 pkpan NMR 64.7%
  • 280802 06/12/09 12:33 auto_snap Target Field and Oven Temps snapshot
  • 280801 06/12/09 12:33 auto_snap Target Cell Temps snapshot
  • 280800 06/12/09 12:33 auto_snap Radiation and Ozone snapshot
  • 280799 06/12/09 12:32 auto_snap Happex Charge Asym snapshot
  • 280798 06/12/09 12:32 auto_snap Hall A Tools snapshot
  • 280797 06/12/09 12:32 auto_snap Compton Polarimeter snapshot
  • 280796 06/12/09 12:32 auto_snap BigBox Magnet snapshot
  • 280795 06/12/09 12:32 auto_snap Beam Position and Current snapshot
  • 280794 06/12/09 12:32 auto_snap Accelerator Energies snapshot
  • 280793 06/12/09 09:17 Higinbotham Right Spectrometer Moved To 17 Degrees
  • 280791 06/12/09 08:34 pkpan NMR 64%
  • 280789 06/12/09 07:58 gilman configuration being changed
  • 280788 06/12/09 07:57 gilman owl shift summary Friday 6/12/09
  • 280785 06/12/09 07:34 gilman 7:30 spot move
  • 280784 06/12/09 07:14 Yawei New happex run started, run#31519
  • 280767 06/12/09 05:17 gilman Hall A tols for shift checklist
  • 280766 06/12/09 05:14 Yawei spot++ check
  • 280761 06/12/09 04:40 Lamiaa Junk runs #: 3506 & 22373 due to 100% dead time problem of R-arm
  • 280760 06/12/09 04:32 auto_snap Target Field and Oven Temps snapshot
  • 280759 06/12/09 04:32 auto_snap Target Cell Temps snapshot
  • 280758 06/12/09 04:32 auto_snap Radiation and Ozone snapshot
  • 280757 06/12/09 04:32 auto_snap Happex Charge Asym snapshot
  • 280756 06/12/09 04:32 auto_snap Hall A Tools snapshot
  • 280755 06/12/09 04:32 auto_snap Compton Polarimeter snapshot
  • 280754 06/12/09 04:32 auto_snap BigBox Magnet snapshot
  • 280753 06/12/09 04:32 auto_snap Beam Position and Current snapshot
  • 280752 06/12/09 04:32 auto_snap Accelerator Energies snapshot
  • 280748 06/12/09 04:26 Yawei NMR 65%
  • 280739 06/12/09 03:18 Lamiaa Restarted new Happex run # 31518
  • 280726 06/12/09 01:55 miham Target Laser Displays still showing errors
  • 280725 06/12/09 01:48 Yawei Laser absorption spectrum after laser trip
  • 280724 06/12/09 01:41 miham Target Lasers tripped by mistake
  • 280719 06/12/09 01:29 miham New golden root files (#22367 and #3500)
  • 280714 06/12/09 00:38 Yawei NMR before taking polarized 3He production run, polarization 66.8%, Laser Absorption Spectrum
  • 280711 06/12/09 00:27 Yawei Target moved to polarized 3He
  • 280707 06/12/09 00:06 Gao Swing shift summary
  • Thursday

  • 280706 06/11/09 23:53 Yi Qiang Spot++ for 3He reference run: 3498
  • 280705 06/11/09 23:49 Yi Qiang Reference Cell Filled with 3He: 135 psig
  • 280696 06/11/09 23:16 miham BigBite sieve slit is OUT
  • 280693 06/11/09 22:59 A. Deur Ref. cell vented
  • 280665 06/11/09 20:35 Yi Qiang Beam current reduced to 5 uA before run 3488/22356
  • 280664 06/11/09 20:32 auto_snap Target Field and Oven Temps snapshot
  • 280663 06/11/09 20:32 auto_snap Target Cell Temps snapshot
  • 280662 06/11/09 20:32 auto_snap Radiation and Ozone snapshot
  • 280661 06/11/09 20:31 auto_snap Happex Charge Asym snapshot
  • 280660 06/11/09 20:31 auto_snap Hall A Tools snapshot
  • 280659 06/11/09 20:31 auto_snap Compton Polarimeter snapshot
  • 280658 06/11/09 20:31 auto_snap BigBox Magnet snapshot
  • 280657 06/11/09 20:31 auto_snap Beam Position and Current snapshot
  • 280656 06/11/09 20:31 auto_snap Accelerator Energies snapshot
  • 280655 06/11/09 20:27 Yi Qiang spot++ check for run 3448
  • 280653 06/11/09 20:20 Yi Qiang Right HRS momentum set to 2.025 GeV/c
  • 280647 06/11/09 19:57 brads HRS sieves OUT, BigBite sieve IN
  • 280646 06/11/09 19:38 Yi Qiang Controlled access to change sieves
  • 280641 06/11/09 18:43 adaq Compton Electron detector in Compton position
  • 280640 06/11/09 18:39 A. Deur Ref cell filled with 135 psig of H2
  • 280625 06/11/09 16:42 V. Sulkosky Beam Update
  • 280624 06/11/09 16:36 A. Camsonne Compton Electron detector in garage for beam tune
  • 280623 06/11/09 16:27 Higinbotham
  • 280621 06/11/09 16:02 Gao Day shift summary
  • 280614 06/11/09 13:33 Yi Qiang H2 bottle connected to the reference cell
  • 280613 06/11/09 13:16 luhj Reference Cell Vacuum
  • 280612 06/11/09 13:15 Higinbotham ND Moved and Sieve Installed
  • 280611 06/11/09 12:52 luhj right arm momemtum has been set to 2.175GeV/c
  • 280610 06/11/09 12:31 auto_snap Target Field and Oven Temps snapshot
  • 280609 06/11/09 12:31 auto_snap Target Cell Temps snapshot
  • 280608 06/11/09 12:31 auto_snap Radiation and Ozone snapshot
  • 280607 06/11/09 12:31 auto_snap Happex Charge Asym snapshot
  • 280606 06/11/09 12:31 auto_snap Hall A Tools snapshot
  • 280605 06/11/09 12:31 auto_snap Compton Polarimeter snapshot
  • 280604 06/11/09 12:31 auto_snap BigBox Magnet snapshot
  • 280603 06/11/09 12:31 auto_snap Beam Position and Current snapshot
  • 280602 06/11/09 12:31 auto_snap Accelerator Energies snapshot
  • 280601 06/11/09 10:56 R. Subedi Re: Summary of Pion-Test-Run Activity
  • 280600 06/11/09 10:49 R. Subedi Summary of Pion-Test-Run Activity
  • 280599 06/11/09 09:38 Higinbotham Planned Configuration Changes
  • 280594 06/11/09 08:48 Luhj Target moved to Reference Cell
  • 280590 06/11/09 08:21 Luhj NMR62.57
  • 280587 06/11/09 08:16 miham Disk Full on adaql10
  • 280586 06/11/09 08:14 jpchen owl shift summary
  • 280585 06/11/09 08:07 miham RTD1 temperatures on Joint N6 during the Night
  • 280579 06/11/09 07:22 pkpan start a new Happex run 31513 at 07:04AM
  • 280578 06/11/09 07:19 pkpan Compton run 17307 still "Run Ended" failing, and do "coda reboot" twice, change the run# to 17308 manually
  • 280577 06/11/09 07:16 X. Zheng Quick analysis for pion trigger efficiencies
  • 280568 06/11/09 06:05 pkpan spot++ snapshot for Run 22343
  • 280559 06/11/09 04:52 R. Subedi Pion signal
  • 280554 06/11/09 04:30 auto_snap Target Field and Oven Temps snapshot
  • 280553 06/11/09 04:30 auto_snap Target Cell Temps snapshot
  • 280552 06/11/09 04:30 auto_snap Radiation and Ozone snapshot
  • 280551 06/11/09 04:30 auto_snap Happex Charge Asym snapshot
  • 280550 06/11/09 04:30 auto_snap Hall A Tools snapshot
  • 280549 06/11/09 04:30 auto_snap Compton Polarimeter snapshot
  • 280548 06/11/09 04:30 auto_snap BigBox Magnet snapshot
  • 280547 06/11/09 04:30 auto_snap Beam Position and Current snapshot
  • 280546 06/11/09 04:30 auto_snap Accelerator Energies snapshot
  • 280545 06/11/09 04:20 R. Subedi Happex Run without VNC, no happex data since 10 PM last night
  • 280542 06/11/09 04:06 pkpan NMR 61.6%
  • 280535 06/11/09 03:29 pkpan Previous Compton Run 17305 "Run Ended" failing.
  • 280534 06/11/09 03:29 pkpan Previous Compton Run 17305 "Run Ended" failing.
  • 280533 06/11/09 03:28 pkpan previous Happex Run 31509 "Run Ended" failing.
  • 280532 06/11/09 03:28 pkpan previous Happex Run 31509 "Run Ended" failing.
  • 280531 06/11/09 03:18 R. Subedi Happex "Run End" failing
  • 280524 06/11/09 02:46 jpchen iochla rebooted.
  • 280523 06/11/09 02:27 jpchen hall a line ioc problem beam down
  • 280519 06/11/09 01:56 pkpan stat a new Happex run 31510.
  • 280518 06/11/09 01:55 pkpan start a new compton run 17306
  • 280512 06/11/09 01:39 jpchen Right p0 is still changing, but close. Started taking data. Note: T3=0.
  • 280504 06/11/09 00:16 Gao Swing shift summary
  • 280503 06/11/09 00:15 jpchen change right HRS to 0.7 GeV
  • Wednesday

  • 280494 06/10/09 23:22 Markowitz NMR 60.4%
  • 280470 06/10/09 20:30 auto_snap Target Field and Oven Temps snapshot
  • 280469 06/10/09 20:30 auto_snap Target Cell Temps snapshot
  • 280468 06/10/09 20:30 auto_snap Radiation and Ozone snapshot
  • 280467 06/10/09 20:30 auto_snap Happex Charge Asym snapshot
  • 280466 06/10/09 20:30 auto_snap Hall A Tools snapshot
  • 280465 06/10/09 20:30 auto_snap Compton Polarimeter snapshot
  • 280464 06/10/09 20:30 auto_snap BigBox Magnet snapshot
  • 280463 06/10/09 20:30 auto_snap Beam Position and Current snapshot
  • 280462 06/10/09 20:30 auto_snap Accelerator Energies snapshot
  • 280455 06/10/09 19:37 X. Qian spot++ saved plot
  • 280452 06/10/09 19:17 Yi/Miha Joint in the end of fiber N6 is hotter
  • 280451 06/10/09 19:06 Markowitz NMR 51.2%
  • 280448 06/10/09 19:05 miham RUTGERS Laser now connected to Fiber N6.
  • 280445 06/10/09 18:59 Yi/Miha Replaced damaged jumper fiber for the comet lasers
  • 280444 06/10/09 18:51 Yi/Miha Replaced damaged jumper fiber for the comet lasers
  • 280443 06/10/09 18:45 Yi Qiang NMR: 51.18%
  • 280440 06/10/09 18:32 Yi/Miha Laser Power Test
  • 280437 06/10/09 18:23 H. Gao Beam half-wave plate is in
  • 280425 06/10/09 16:15 Markowitz NMR 60%
  • 280419 06/10/09 16:04 averett day shift summary
  • 280414 06/10/09 15:25 averett Wavetek 80 fan noise
  • 280409 06/10/09 14:55 averett Hall A tools screen
  • 280408 06/10/09 14:54 averett spot ++
  • 280406 06/10/09 14:32 miham Quick analysis of N2 runs taken last night
  • 280402 06/10/09 14:17 miham Target Polarization and Field Orientation Plots
  • 280393 06/10/09 12:29 auto_snap Target Field and Oven Temps snapshot
  • 280392 06/10/09 12:29 auto_snap Target Cell Temps snapshot
  • 280391 06/10/09 12:29 auto_snap Radiation and Ozone snapshot
  • 280390 06/10/09 12:29 auto_snap Happex Charge Asym snapshot
  • 280389 06/10/09 12:29 auto_snap Hall A Tools snapshot
  • 280388 06/10/09 12:29 auto_snap Compton Polarimeter snapshot
  • 280387 06/10/09 12:29 auto_snap BigBox Magnet snapshot
  • 280386 06/10/09 12:29 auto_snap Beam Position and Current snapshot
  • 280385 06/10/09 12:29 auto_snap Accelerator Energies snapshot
  • 280382 06/10/09 12:13 Yi Qiang Target Spin Rotated to Longitudinal+ (0 degree)
  • 280371 06/10/09 10:29 adaq spot++ for run 3440
  • 280368 06/10/09 10:24 adaq spot++ for run 3439
  • 280365 06/10/09 10:16 gjin half-wave plate out for this run and after
  • 280364 06/10/09 10:16 gjin have-wave plate out for this run and the runs after
  • 280357 06/10/09 08:57 averett BHWP is now OUT
  • 280356 06/10/09 08:53 averett NMR=58.2%
  • 280348 06/10/09 07:59 ejensen Owl shift summary
  • 280335 06/10/09 06:15 ejensen SPOT++ check
  • 280334 06/10/09 06:14 ejensen Taking beam
  • 280331 06/10/09 05:08 ejensen Beam down
  • 280327 06/10/09 04:47 Markowitz NMR 56.7
  • 280319 06/10/09 04:29 auto_snap Target Field and Oven Temps snapshot
  • 280318 06/10/09 04:29 auto_snap Target Cell Temps snapshot
  • 280317 06/10/09 04:29 auto_snap Radiation and Ozone snapshot
  • 280316 06/10/09 04:29 auto_snap Happex Charge Asym snapshot
  • 280315 06/10/09 04:29 auto_snap Hall A Tools snapshot
  • 280314 06/10/09 04:29 auto_snap Compton Polarimeter snapshot
  • 280313 06/10/09 04:29 auto_snap BigBox Magnet snapshot
  • 280312 06/10/09 04:29 auto_snap Beam Position and Current snapshot
  • 280311 06/10/09 04:28 auto_snap Accelerator Energies snapshot
  • 280310 06/10/09 04:27 ejensen SPOT++ check
  • 280307 06/10/09 04:05 ejensen Beam trip shorter than expected
  • 280304 06/10/09 04:01 ejensen Beam down temporarily
  • 280303 06/10/09 03:26 Markowitz NMR 58.8%
  • 280298 06/10/09 03:18 Lamiaa End run #22298 is missing due to end run failed
  • 280297 06/10/09 03:03 miham New golden root file for HRS-R and ND: #22298
  • 280293 06/10/09 02:26 Lamiaa rightarm_p=1.865 insted of rightarm_p=2.175
  • 280292 06/10/09 02:24 ejensen Started new compton run #17289
  • 280291 06/10/09 02:23 ejensen Started new HAPPEX run #31504
  • 280290 06/10/09 02:22 ejensen Begin normal production with runs 3427 and 22298*
  • 280289 06/10/09 02:17 ejensen Right HMS momentum stabilized at 1.865 GeV/c
  • 280276 06/10/09 00:43 miham Target Strip charts
  • 280274 06/10/09 00:40 miham NMR frequency sweep (polarization 57.98% )
  • 280272 06/10/09 00:38 S. Sirca Start ramp-down of Right HRS momentum to 1.865 GeV/c
  • 280271 06/10/09 00:35 miham Target moved to 3He cell
  • 280270 06/10/09 00:32 miham Reference cell filled with N2 for Calibration runs
  • 280266 06/10/09 00:29 S. Sirca Raw rates, ref.cell N2, 18 psig, run #3423, 9 uA
  • 280263 06/10/09 00:27 pkpan swing shift summary / June 9
  • 280260 06/10/09 00:24 S. Sirca Raw rates, ref. cell N2, 15 psig, run #3422, 9 uA
  • 280255 06/10/09 00:18 S. Sirca Raw rates, ref.cell N2, 12 psig, run #3421, 9 uA
  • 280250 06/10/09 00:15 miham Run #22290 is a calibration run with 6psig of N2 in reference cell
  • 280249 06/10/09 00:15 S. Sirca Raw rates, ref.cell N2, 9 psig, run #3420, 9 uA
  • 280244 06/10/09 00:10 S. Sirca Raw rates, ref.cell N2, 6 psig, run #3419, 9 uA
  • 280239 06/10/09 00:03 S. Sirca Raw rates, ref.cell 3 psig, run #3418, 9 uA
  • Tuesday

  • 280229 06/09/09 23:48 S. Sirca Raw rates, ref.cell N2, 0 psig, run #3415, 9 uA
  • 280224 06/09/09 23:42 S. Sirca Raw rates, ref.cell N2 -3psi, run #3414, 9 uA
  • 280219 06/09/09 23:31 S. Sirca Raw rates, ref.cell N2 -6 psig, run #3413, 9 uA
  • 280214 06/09/09 23:25 S. Sirca Raw rates, ref.cell N2 -9 psig, run #3412, 9 uA
  • 280209 06/09/09 23:20 S. Sirca Raw rates, ref.cell N2, -12psig, run #3411, 9uA
  • 280204 06/09/09 23:11 S. Sirca Raw rates, ref.cell empty, run #3410
  • 280201 06/09/09 22:57 pkpan End run of 3407 is missing
  • 280200 06/09/09 22:45 X. Zheng target moved to reference cell
  • 280197 06/09/09 22:13 pkpan start Happex run 31503
  • 280192 06/09/09 22:06 brads Controlled access from 21:20-21:35 to pull BB Sieve
  • 280191 06/09/09 22:04 adaq spot++ for 3He run 3408, raster 4X4
  • 280190 06/09/09 21:57 Bo Run 22276 is a production run
  • 280184 06/09/09 21:34 brads BigBite Sieve is OUT
  • 280175 06/09/09 20:33 pkpan NMR 59.1%
  • 280172 06/09/09 20:28 auto_snap Target Field and Oven Temps snapshot
  • 280171 06/09/09 20:28 auto_snap Target Cell Temps snapshot
  • 280170 06/09/09 20:28 auto_snap Radiation and Ozone snapshot
  • 280169 06/09/09 20:28 auto_snap Happex Charge Asym snapshot
  • 280168 06/09/09 20:28 auto_snap Hall A Tools snapshot
  • 280167 06/09/09 20:28 auto_snap Compton Polarimeter snapshot
  • 280166 06/09/09 20:28 auto_snap BigBox Magnet snapshot
  • 280165 06/09/09 20:28 auto_snap Beam Position and Current snapshot
  • 280164 06/09/09 20:28 auto_snap Accelerator Energies snapshot
  • 280159 06/09/09 20:01 adaq spot++ for 3He, raster 4X4
  • 280150 06/09/09 19:38 adaq spot++ for 3He, raster 4X4
  • 280145 06/09/09 19:18 pkpan spot++ for carbon, raster 4X4
  • 280140 06/09/09 17:45 miham Spot++ check during calibration run #3399
  • 280139 06/09/09 17:43 miham Runs #3398 and #22270 - RASTER OFF
  • 280129 06/09/09 15:59 Annand Day shift summary
  • 280122 06/09/09 15:12 Yi Qiang Reference cell filled with D2: 135 psig
  • 280121 06/09/09 15:02 Yi Qiang A "new" D2 bottle has been connected to the reference cell
  • 280120 06/09/09 14:34 Bo Beam status update
  • 280119 06/09/09 12:28 Yi Qiang Q-wave plates tests for transverse pumping
  • 280118 06/09/09 12:28 auto_snap Target Field and Oven Temps snapshot
  • 280117 06/09/09 12:28 auto_snap Target Cell Temps snapshot
  • 280116 06/09/09 12:28 Yi Qiang Q-wave plates tests for transverse pumping
  • 280115 06/09/09 12:28 auto_snap Radiation and Ozone snapshot
  • 280114 06/09/09 12:28 auto_snap Happex Charge Asym snapshot
  • 280113 06/09/09 12:28 auto_snap Hall A Tools snapshot
  • 280112 06/09/09 12:28 auto_snap Compton Polarimeter snapshot
  • 280111 06/09/09 12:28 auto_snap BigBox Magnet snapshot
  • 280110 06/09/09 12:28 auto_snap Beam Position and Current snapshot
  • 280109 06/09/09 12:27 auto_snap Accelerator Energies snapshot
  • 280108 06/09/09 12:22 Yi Qiang NMR After Internal Temperature Test: 49%
  • 280107 06/09/09 12:15 Yi Qiang D2 bottle is connected to reference cell, but it's almost empty
  • 280106 06/09/09 12:02 Yi Qiang Internal Temperature Test for Transverse Pumping
  • 280105 06/09/09 11:26 R. Michaels possible clock asymmetry
  • 280104 06/09/09 10:24 Higinbotham BigBite Sieve Installed
  • 280103 06/09/09 10:18 Higinbotham Right HRS Quads Cycled
  • 280102 06/09/09 10:18 Yi Qiang NMR: 59.19%
  • 280101 06/09/09 10:14 Higinbotham HRS Front and Back Floor Marks [a.k.a. poor man's survey]
  • 280100 06/09/09 10:02 Ramesh/Xiaoc Access activity [pvdis work]
  • 280099 06/09/09 08:50 Bo Hall is in Controlled Access
  • 280098 06/09/09 08:27 Annand No hard estimate of beam studies period
  • 280096 06/09/09 08:00 ejensen Owl shift summary
  • 280095 06/09/09 07:16 ejensen Beam studies
  • 280088 06/09/09 06:06 Lamiaa Restarted new Happex/Compton runs @ 6:00 AM
  • 280085 06/09/09 05:56 Markowitz NMR 58.05%
  • 280074 06/09/09 04:27 auto_snap Target Field and Oven Temps snapshot
  • 280073 06/09/09 04:27 auto_snap Target Cell Temps snapshot
  • 280072 06/09/09 04:27 auto_snap Radiation and Ozone snapshot
  • 280071 06/09/09 04:27 auto_snap Happex Charge Asym snapshot
  • 280070 06/09/09 04:27 auto_snap Hall A Tools snapshot
  • 280069 06/09/09 04:27 auto_snap Compton Polarimeter snapshot
  • 280068 06/09/09 04:27 auto_snap BigBox Magnet snapshot
  • 280067 06/09/09 04:27 auto_snap Beam Position and Current snapshot
  • 280066 06/09/09 04:27 auto_snap Accelerator Energies snapshot
  • 280061 06/09/09 03:45 ejensen New compton run
  • 280052 06/09/09 03:03 ejensen SPOT++ check
  • 280051 06/09/09 02:46 Lamiaa Restarted new Happex/Compton runs @ 2:00 AM
  • 280050 06/09/09 02:46 Lamiaa Restarted new Happex/Compton runs @ 2:00 AM
  • 280045 06/09/09 02:05 Markowitz absorption spectrum
  • 280043 06/09/09 02:03 Markowitz NMR 58.5%
  • 280026 06/09/09 00:00 S. Sirca Excessive (?) beam position oscillations
  • Monday

  • 280024 06/08/09 23:57 adaq Swing shift SUMMARY
  • 280008 06/08/09 21:55 miham Target strip charts after target moved to 3He
  • 280007 06/08/09 21:52 deng NMR 56.98%
  • 280004 06/08/09 21:50 S. Sirca New Compton run #17265
  • 280003 06/08/09 21:47 deng NMR 56.98%
  • 280002 06/08/09 21:44 miham Target moved to 3He Position
  • 279999 06/08/09 21:31 S. Sirca Dead times ref.cell N2 vs. ref.cell H2
  • 279998 06/08/09 21:31 miham Reference cell filled with N2
  • 279997 06/08/09 21:28 S. Sirca Raw rates ref.cell N2, 9uA, run #3378
  • 279984 06/08/09 20:27 auto_snap Target Field and Oven Temps snapshot
  • 279983 06/08/09 20:27 auto_snap Target Cell Temps snapshot
  • 279982 06/08/09 20:27 auto_snap Radiation and Ozone snapshot
  • 279981 06/08/09 20:27 auto_snap Happex Charge Asym snapshot
  • 279980 06/08/09 20:27 auto_snap Hall A Tools snapshot
  • 279979 06/08/09 20:27 auto_snap Compton Polarimeter snapshot
  • 279978 06/08/09 20:27 auto_snap BigBox Magnet snapshot
  • 279977 06/08/09 20:26 auto_snap Beam Position and Current snapshot
  • 279976 06/08/09 20:26 auto_snap Accelerator Energies snapshot
  • 279971 06/08/09 19:59 miham Spot++ check during Hydrogen run #3374
  • 279970 06/08/09 19:57 S. Sirca Raw rates ref.cell H2, 9uA, T5 only
  • 279965 06/08/09 19:47 miham Spot++ check during Hydrogen run #3373
  • 279964 06/08/09 19:45 S. Sirca Raw rates, ref.cell H2, 9uA, BB singles ONLY
  • 279957 06/08/09 19:32 miham Target moved to Reference Cell
  • 279954 06/08/09 19:15 miham Spot++ check during Optics run #3371 - Raster off
  • 279953 06/08/09 19:11 S. Sirca Raw rates with Optics/BeO 5uA run #3371, raster OFF
  • 279948 06/08/09 19:03 miham Spot++ check during Optics Run #3370
  • 279947 06/08/09 19:03 S. Sirca y-target sanity check run #3370
  • 279946 06/08/09 19:00 S. Sirca Raw rates with Optics/BeO 5uA run #3370, raster ON
  • 279945 06/08/09 18:54 S. Sirca New Compton run #17264
  • 279938 06/08/09 18:41 miham Target Moved to Optics
  • 279935 06/08/09 18:23 deng screen shot for laser absorption spectrum
  • 279934 06/08/09 18:18 miham Laser problem fixed
  • 279929 06/08/09 18:02 miham Target laser 1 apparently tripped
  • 279928 06/08/09 17:42 miham Beam jiggling around
  • 279915 06/08/09 16:05 Annand Day shift summary
  • 279913 06/08/09 15:56 Bo Start new compton run 17262
  • 279904 06/08/09 14:55 Annand Compton new run 17261....missing runs? Remote interference?
  • 279901 06/08/09 14:30 De Leo NMR Control
  • 279898 06/08/09 14:13 Annand Significant accelerator down time
  • 279893 06/08/09 13:13 Annand New Happex run 31498 at 12.51
  • 279892 06/08/09 13:09 Annand spot++
  • 279887 06/08/09 12:43 Annand New Compton run 17257 (end run 17255 failed, 17256 JUNK)
  • 279886 06/08/09 12:26 auto_snap Target Field and Oven Temps snapshot
  • 279885 06/08/09 12:26 auto_snap Target Cell Temps snapshot
  • 279884 06/08/09 12:26 auto_snap Radiation and Ozone snapshot
  • 279883 06/08/09 12:26 auto_snap Happex Charge Asym snapshot
  • 279882 06/08/09 12:26 auto_snap Hall A Tools snapshot
  • 279881 06/08/09 12:26 auto_snap Compton Polarimeter snapshot
  • 279880 06/08/09 12:26 auto_snap BigBox Magnet snapshot
  • 279879 06/08/09 12:26 auto_snap Beam Position and Current snapshot
  • 279878 06/08/09 12:26 auto_snap Accelerator Energies snapshot
  • 279867 06/08/09 11:05 Annand Lost beam ~20 mins
  • 279864 06/08/09 10:36 Annand New Compton run 17255 (end run 17254 failed)
  • 279859 06/08/09 10:02 De Leo NMR 57.94
  • 279848 06/08/09 08:47 Annand New Happex run 31497
  • 279845 06/08/09 08:37 Annand New Compton run 17254
  • 279839 06/08/09 08:02 Markowitz Owl shift summary
  • 279838 06/08/09 07:36 ejensen ROC 5 crash (runs 3348 3349 22221 22222)
  • 279819 06/08/09 05:28 Markowitz Tools screenshot
  • 279817 06/08/09 05:27 Markowitz NMR 57.55%
  • 279813 06/08/09 05:23 Markowitz Spot++ screenshot
  • 279812 06/08/09 05:15 miham dE/E plot for elastic run #3265
  • 279807 06/08/09 04:26 auto_snap Target Field and Oven Temps snapshot
  • 279806 06/08/09 04:26 auto_snap Target Cell Temps snapshot
  • 279805 06/08/09 04:26 auto_snap Radiation and Ozone snapshot
  • 279804 06/08/09 04:26 auto_snap Happex Charge Asym snapshot
  • 279803 06/08/09 04:26 auto_snap Hall A Tools snapshot
  • 279802 06/08/09 04:26 auto_snap Compton Polarimeter snapshot
  • 279801 06/08/09 04:26 auto_snap BigBox Magnet snapshot
  • 279800 06/08/09 04:25 auto_snap Beam Position and Current snapshot
  • 279799 06/08/09 04:25 auto_snap Accelerator Energies snapshot
  • 279790 06/08/09 03:36 miham Hydrogen elastic run #3265 analysis
  • 279789 06/08/09 03:28 ejensen CODA issues (runs 3339, 3340, 22212, 22213)
  • 279770 06/08/09 01:38 Markowitz absorption spectrum
  • 279768 06/08/09 01:34 Markowitz NMR 56.75%
  • 279756 06/08/09 00:01 S. Sirca Swing shift SUMMARY
  • 279755 06/08/09 00:01 LeRose absorption spectrum
  • Sunday

  • 279751 06/07/09 23:45 S. Sirca Beam OFF
  • 279746 06/07/09 23:38 S. Sirca New HAPPEX run #31495 started
  • 279728 06/07/09 21:31 S. Sirca New Compton run #17249
  • 279727 06/07/09 21:24 LeRose NMR really 57.8994%
  • 279726 06/07/09 21:22 LeRose NMR 57.1%
  • 279717 06/07/09 20:25 auto_snap Target Field and Oven Temps snapshot
  • 279716 06/07/09 20:25 auto_snap Target Cell Temps snapshot
  • 279715 06/07/09 20:25 auto_snap Radiation and Ozone snapshot
  • 279714 06/07/09 20:25 auto_snap Happex Charge Asym snapshot
  • 279713 06/07/09 20:25 auto_snap Hall A Tools snapshot
  • 279712 06/07/09 20:25 auto_snap Compton Polarimeter snapshot
  • 279711 06/07/09 20:25 auto_snap BigBox Magnet snapshot
  • 279710 06/07/09 20:25 auto_snap Beam Position and Current snapshot
  • 279709 06/07/09 20:25 auto_snap Accelerator Energies snapshot
  • 279704 06/07/09 20:15 S. Sirca First truly good Compton run after HWP=IN is #17248
  • 279703 06/07/09 20:13 A. Camsonne A. Camsonne
  • 279698 06/07/09 19:25 A. Camsonne Stand alone Compton Electron detector photon electon coincidence Run 17245
  • 279695 06/07/09 19:14 S. Sirca Half-wave plate (HWP) now IN
  • 279684 06/07/09 17:25 LeRose NMR Sweep (57.4%)
  • 279675 06/07/09 16:35 S. Sirca Compton run #17234, HV trip, run #17235
  • 279670 06/07/09 16:02 Annand Day shift summary
  • 279668 06/07/09 15:54 Yi Qiang Q-wave plates tests for transverse pumping
  • 279655 06/07/09 14:19 Annand Compton new run 17233
  • 279648 06/07/09 13:36 De Leo NMR Frequency control
  • 279639 06/07/09 12:32 Annand New Happex run 31492
  • 279636 06/07/09 12:25 auto_snap Target Field and Oven Temps snapshot
  • 279635 06/07/09 12:25 auto_snap Target Cell Temps snapshot
  • 279634 06/07/09 12:25 auto_snap Radiation and Ozone snapshot
  • 279633 06/07/09 12:25 auto_snap Happex Charge Asym snapshot
  • 279632 06/07/09 12:25 auto_snap Hall A Tools snapshot
  • 279631 06/07/09 12:25 auto_snap Compton Polarimeter snapshot
  • 279630 06/07/09 12:24 auto_snap BigBox Magnet snapshot
  • 279629 06/07/09 12:24 auto_snap Beam Position and Current snapshot
  • 279628 06/07/09 12:24 auto_snap Accelerator Energies snapshot
  • 279627 06/07/09 12:19 Annand New Compton run 17232
  • 279626 06/07/09 12:00 Annand spot++
  • 279613 06/07/09 10:11 Annand New Compton run 17231
  • 279608 06/07/09 09:36 Annand Junk 3306, 22178
  • 279607 06/07/09 09:35 Annand Restart RH coda: run 22178 failed
  • 279604 06/07/09 09:29 De Leo NMR Frequency control
  • 279598 06/07/09 08:50 Annand Run 3304 Junk
  • 279591 06/07/09 08:18 Annand New Happex run 31491
  • 279586 06/07/09 08:08 miham New Compton Run #17230
  • 279585 06/07/09 08:05 miham OWL Shift Summary (06/07/09)
  • 279579 06/07/09 07:32 pkpan Laser absorption spectrum 06_07_0730
  • 279574 06/07/09 06:57 miham Bad beam for the last hour.
  • 279565 06/07/09 06:11 miham/Kai New Compton run #17229
  • 279562 06/07/09 05:38 pkpan NMR 57.1%
  • 279551 06/07/09 04:24 auto_snap Target Field and Oven Temps snapshot
  • 279550 06/07/09 04:24 auto_snap Target Cell Temps snapshot
  • 279549 06/07/09 04:24 auto_snap Radiation and Ozone snapshot
  • 279548 06/07/09 04:24 auto_snap Happex Charge Asym snapshot
  • 279547 06/07/09 04:24 auto_snap Hall A Tools snapshot
  • 279546 06/07/09 04:24 auto_snap Compton Polarimeter snapshot
  • 279545 06/07/09 04:24 auto_snap BigBox Magnet snapshot
  • 279544 06/07/09 04:24 auto_snap Beam Position and Current snapshot
  • 279543 06/07/09 04:24 auto_snap Accelerator Energies snapshot
  • 279538 06/07/09 03:52 pkpan New compton run #17228.
  • 279521 06/07/09 01:49 miham New Compton run #17227 after ROC reboot
  • 279520 06/07/09 01:35 miham BB, ND, HRS-R HV Gui screen shot - All green.
  • 279519 06/07/09 01:31 mihan Spot++ check (#3290)
  • 279516 06/07/09 01:25 pkpan NMR 57.5%
  • 279513 06/07/09 00:59 miham ROC 5 problem during prod. run #3288
  • 279512 06/07/09 00:56 miham Changed the beam current back to 9uA (original production setting)
  • 279504 06/07/09 00:22 miham Beam current lowered to 5uA for runs #3287/#22160
  • 279499 06/07/09 00:08 S. Sirca Swing shift SUMMARY
  • Saturday

  • 279493 06/06/09 23:25 S. Sirca New Compton run #17225
  • 279480 06/06/09 21:45 Shahinyan NMR_06_Jun_09_2
  • 279471 06/06/09 20:59 adaq Compton DAQ died, restarted, new run #17224
  • 279466 06/06/09 20:30 S New HAPPEX run #31488 started
  • 279465 06/06/09 20:24 auto_snap Target Field and Oven Temps snapshot
  • 279464 06/06/09 20:24 auto_snap Target Cell Temps snapshot
  • 279463 06/06/09 20:23 auto_snap Radiation and Ozone snapshot
  • 279462 06/06/09 20:23 auto_snap Happex Charge Asym snapshot
  • 279461 06/06/09 20:23 auto_snap Hall A Tools snapshot
  • 279460 06/06/09 20:23 auto_snap Compton Polarimeter snapshot
  • 279459 06/06/09 20:23 auto_snap BigBox Magnet snapshot
  • 279458 06/06/09 20:23 auto_snap Beam Position and Current snapshot
  • 279457 06/06/09 20:23 auto_snap Accelerator Energies snapshot
  • 279448 06/06/09 18:58 S. Sirca Compton run #17222 started
  • 279443 06/06/09 18:37 shahinyan target
  • 279438 06/06/09 17:55 Shahinyan NMR_6_jun_2009
  • 279425 06/06/09 16:18 S. Sirca Compton run #17221 started
  • 279420 06/06/09 16:01 Annand Day shift summary
  • 279417 06/06/09 15:48 Annand New Happex run 31487
  • 279397 06/06/09 13:38 Annand spot++
  • 279396 06/06/09 13:38 Bo Set new golden run 3266 for L-arm and 22139 for R-arm
  • 279394 06/06/09 13:07 Target spin rotated to Transverse - (270 degrees)
  • 279390 06/06/09 12:31 A. Camsonne Compton electron detector checks
  • 279387 06/06/09 12:26 Annand Move to ref cell and take data
  • 279386 06/06/09 12:25 A. Camsonne Compton electron detector in
  • 279385 06/06/09 12:25 Yi Qiang Reference cell filled with H2: 135 psig
  • 279384 06/06/09 12:23 auto_snap Target Field and Oven Temps snapshot
  • 279383 06/06/09 12:23 auto_snap Target Cell Temps snapshot
  • 279382 06/06/09 12:23 auto_snap Radiation and Ozone snapshot
  • 279381 06/06/09 12:23 auto_snap Happex Charge Asym snapshot
  • 279380 06/06/09 12:23 auto_snap Hall A Tools snapshot
  • 279379 06/06/09 12:23 auto_snap Compton Polarimeter snapshot
  • 279378 06/06/09 12:23 auto_snap BigBox Magnet snapshot
  • 279377 06/06/09 12:23 auto_snap Beam Position and Current snapshot
  • 279376 06/06/09 12:23 auto_snap Accelerator Energies snapshot
  • 279372 06/06/09 11:52 Annand Happex new run 31486
  • 279364 06/06/09 10:53 DeLeo NMR frequency control
  • 279343 06/06/09 08:16 Annand Happex restart 08.06
  • 279340 06/06/09 08:03 Ibrahim Owl Shift Summary (June 6, 2009)
  • 279339 06/06/09 08:03 Ibrahim No Beam at the End of the Owl Shift
  • 279337 06/06/09 08:00 Lamiaa No beam in the last runs 3256 & 22129
  • 279330 06/06/09 07:19 Ibrahim Restarted the HAPPEX DAQ at 7:06 am. New Run 31483
  • 279329 06/06/09 07:19 Lamiaa Spot++ w/. 4x4 mm raster after beam is back
  • 279328 06/06/09 07:19 Ibrahim No beam for 15 min.
  • 279327 06/06/09 07:10 Lamiaa Runs 3254/22127 maybe junk!
  • 279316 06/06/09 06:14 pkpan NMR 61.9%
  • 279313 06/06/09 05:44 Ibrahim Restarted the HAPPEX DAQ at 5:41 am. New Run 31482
  • 279300 06/06/09 04:30 Lamiaa Junk runs; LHRS/RHRS: 3247 & 3248/ 22120 & 22121
  • 279299 06/06/09 04:23 auto_snap Target Field and Oven Temps snapshot
  • 279298 06/06/09 04:22 auto_snap Target Cell Temps snapshot
  • 279297 06/06/09 04:22 auto_snap Radiation and Ozone snapshot
  • 279296 06/06/09 04:22 auto_snap Happex Charge Asym snapshot
  • 279295 06/06/09 04:22 auto_snap Hall A Tools snapshot
  • 279294 06/06/09 04:22 auto_snap Compton Polarimeter snapshot
  • 279293 06/06/09 04:22 auto_snap BigBox Magnet snapshot
  • 279292 06/06/09 04:22 auto_snap Beam Position and Current snapshot
  • 279291 06/06/09 04:22 auto_snap Accelerator Energies snapshot
  • 279276 06/06/09 03:26 Ibrahim Started the HAPPEX DAQ at 1:57 am. New Run is 31481
  • 279270 06/06/09 02:31 Ibrahim Spot++ for Run 3243
  • 279268 06/06/09 02:30 pkpan NMR 62.5%
  • 279265 06/06/09 02:19 Ibrahim Beam Chart
  • 279260 06/06/09 01:54 Ibrahim Rates for He3 4x4 mm and 5 uA
  • 279259 06/06/09 01:50 Ibrahim Rates for He3 4x4 mm 2 uA
  • 279256 06/06/09 01:48 pkpan target moved to 3He
  • 279255 06/06/09 01:44 Lamiaa Spot++
  • 279246 06/06/09 01:01 pkpan target moved to the oprtics
  • 279245 06/06/09 00:03 S. Sirca Swing shift SUMMARY
  • 279244 06/06/09 00:02 S. Sirca Compton DAQ recovered
  • Friday

  • 279240 06/05/09 23:31 S. Sirca South Linac RF problems
  • 279231 06/05/09 22:22 Jin Huang NMR 61.1%; NMR code update
  • 279222 06/05/09 21:36 Annand NMR last sweep screen shot not saved properly
  • 279221 06/05/09 21:32 Annand NMR 61.31%
  • 279212 06/05/09 20:31 Annand Laser absorption spectrum
  • 279211 06/05/09 20:30 S. Sirca Run #22105 has no start-of-run entry && can not be replayed
  • 279206 06/05/09 20:22 auto_snap Target Field and Oven Temps snapshot
  • 279205 06/05/09 20:22 auto_snap Target Cell Temps snapshot
  • 279204 06/05/09 20:22 auto_snap Radiation and Ozone snapshot
  • 279203 06/05/09 20:22 auto_snap Happex Charge Asym snapshot
  • 279202 06/05/09 20:22 auto_snap Hall A Tools snapshot
  • 279201 06/05/09 20:22 auto_snap Compton Polarimeter snapshot
  • 279200 06/05/09 20:22 auto_snap BigBox Magnet snapshot
  • 279199 06/05/09 20:22 auto_snap Beam Position and Current snapshot
  • 279198 06/05/09 20:22 auto_snap Accelerator Energies snapshot
  • 279189 06/05/09 19:14 S. Sirca Compton PC dead
  • 279174 06/05/09 17:16 A. Camsonne Summary Compton photon asymmetry
  • 279173 06/05/09 17:10 Annand NMR = 63.29%
  • 279170 06/05/09 17:06 miham HAPPEX Crashed - Probably not working for more than 3 hours.
  • 279164 06/05/09 15:59 gjin day shift summary
  • 279141 06/05/09 12:59 De Leo NMR Frequency Control
  • 279136 06/05/09 12:21 auto_snap Target Field and Oven Temps snapshot
  • 279135 06/05/09 12:21 auto_snap Target Cell Temps snapshot
  • 279134 06/05/09 12:21 auto_snap Radiation and Ozone snapshot
  • 279133 06/05/09 12:21 auto_snap Happex Charge Asym snapshot
  • 279132 06/05/09 12:21 auto_snap Hall A Tools snapshot
  • 279131 06/05/09 12:21 auto_snap Compton Polarimeter snapshot
  • 279130 06/05/09 12:21 auto_snap BigBox Magnet snapshot
  • 279129 06/05/09 12:21 auto_snap Beam Position and Current snapshot
  • 279128 06/05/09 12:21 auto_snap Accelerator Energies snapshot
  • 279105 06/05/09 09:33 gjin halla tools screen
  • 279104 06/05/09 09:32 gjin spot++
  • 279101 06/05/09 09:06 gjin beam down since 820am and now back again
  • 279100 06/05/09 08:45 De Leo NMR Frequency Control
  • 279097 06/05/09 08:39 gjin start new Compton 17206
  • 279096 06/05/09 08:34 gjin cavity problem from MCC. beam back later
  • 279095 06/05/09 08:29 Bo Re:1/2 wave plate entries wrong?
  • 279089 06/05/09 08:01 Holmstrom Owl Shift Summary
  • 279088 06/05/09 07:44 Holmstrom 1/2 wave plate entries wrong?
  • 279071 06/05/09 05:48 adaq Spot++ result
  • 279070 06/05/09 05:44 Holmstrom Hall A tools
  • 279058 06/05/09 04:21 auto_snap Accelerator Energies snapshot
  • 279054 06/05/09 04:11 pkpan NMR 62.2%
  • 279031 06/05/09 00:35 Jin Huang NMR 60.5%
  • 279022 06/05/09 00:05 S. Sirca Swing shift SUMMARY
  • Thursday

  • 279016 06/04/09 23:41 S. Sirca Beam current and position fluctuations
  • 279008 06/04/09 22:42 miham New Golden Root Files : #3193 and #22065
  • 278990 06/04/09 20:21 auto_snap Target Field and Oven Temps snapshot
  • 278989 06/04/09 20:20 auto_snap Target Cell Temps snapshot
  • 278988 06/04/09 20:20 auto_snap Radiation and Ozone snapshot
  • 278987 06/04/09 20:20 auto_snap Happex Charge Asym snapshot
  • 278986 06/04/09 20:20 auto_snap Hall A Tools snapshot
  • 278985 06/04/09 20:20 auto_snap Compton Polarimeter snapshot
  • 278984 06/04/09 20:20 auto_snap BigBox Magnet snapshot
  • 278983 06/04/09 20:20 auto_snap Beam Position and Current snapshot
  • 278982 06/04/09 20:20 auto_snap Accelerator Energies snapshot
  • 278973 06/04/09 19:15 Annand Laser absorption spectrum....with screen shot this time
  • 278972 06/04/09 19:12 S. Sirca Back at 9uA
  • 278971 06/04/09 19:10 Annand Laser absorption spectrum
  • 278970 06/04/09 19:07 Annand NMR 62.64%
  • 278969 06/04/09 19:03 S. Sirca Beam current drops steadily, then abruptly
  • 278968 06/04/09 19:00 S. Sirca New Compton run #17201, new HAPPEX run #31472
  • 278959 06/04/09 18:12 S. Sirca HAND N1 Right TDC0 different than in golden run
  • 278958 06/04/09 17:55 S. Sirca Raw rates with 9uA on He3 at re-start of production
  • 278956 06/04/09 17:53 S. Sirca spot++ with 9uA on pol.He3, 7x5 MCC raster
  • 278949 06/04/09 17:44 S. Sirca 2uA, 6x6 raster on polarized He3
  • 278943 06/04/09 17:34 S. Sirca spot++ check with BeO after beam restored
  • 278938 06/04/09 17:19 miham Target Lasers turned OFF/ON
  • 278937 06/04/09 16:58 S. Sirca Compton HV trips when beam restored
  • 278936 06/04/09 16:18 Annand Target position change to empty
  • 278934 06/04/09 16:00 ellie Day Shift Summary
  • 278933 06/04/09 15:17 A. Camsonne Compton electron detector in garage position
  • 278932 06/04/09 12:20 auto_snap Target Field and Oven Temps snapshot
  • 278931 06/04/09 12:20 auto_snap Target Cell Temps snapshot
  • 278930 06/04/09 12:20 auto_snap Radiation and Ozone snapshot
  • 278929 06/04/09 12:20 auto_snap Happex Charge Asym snapshot
  • 278928 06/04/09 12:20 auto_snap Hall A Tools snapshot
  • 278927 06/04/09 12:20 auto_snap Compton Polarimeter snapshot
  • 278926 06/04/09 12:20 auto_snap BigBox Magnet snapshot
  • 278925 06/04/09 12:20 auto_snap Beam Position and Current snapshot
  • 278924 06/04/09 12:19 auto_snap Accelerator Energies snapshot
  • 278923 06/04/09 12:06 Yi Qiang NMR: 56%
  • 278922 06/04/09 12:05 Yi Qiang NMR Cross Calibration for Transverse + Field
  • 278921 06/04/09 10:53 Ramesh/Xiaoc RHRS shower signals ADC16 and ADC55 back
  • 278920 06/04/09 10:36 Yi Qiang H2 bottle connected to the reference cell (D2 bottle disconnected), Pressured Gauge Zeroed
  • 278917 06/04/09 09:13 Bo Survey is completed and Hall is in controlled access
  • 278916 06/04/09 09:09 ellie Survey Complete, Hall now in Controlled Access
  • 278915 06/04/09 08:16 ellie Compton Having Problems Ending Run
  • 278914 06/04/09 08:12 ellie Ended Compton Run since we're going to Controlled Access
  • 278913 06/04/09 08:11 ellie Ended HAPPEX run since we're going to Controlled Access
  • 278912 06/04/09 08:03 W. Boeglin Shift Summary
  • 278910 06/04/09 07:06 Boeglin Beam studies begin
  • 278905 06/04/09 06:27 Boeglin Compton HV trip
  • 278898 06/04/09 05:38 pkpan NMR 62.9%
  • 278885 06/04/09 04:27 ejensen SPOT++ check
  • 278884 06/04/09 04:19 auto_snap Target Field and Oven Temps snapshot
  • 278883 06/04/09 04:19 auto_snap Target Cell Temps snapshot
  • 278882 06/04/09 04:19 auto_snap Radiation and Ozone snapshot
  • 278881 06/04/09 04:19 auto_snap Happex Charge Asym snapshot
  • 278880 06/04/09 04:19 auto_snap Hall A Tools snapshot
  • 278879 06/04/09 04:19 auto_snap Compton Polarimeter snapshot
  • 278878 06/04/09 04:19 auto_snap BigBox Magnet snapshot
  • 278877 06/04/09 04:19 auto_snap Beam Position and Current snapshot
  • 278876 06/04/09 04:19 auto_snap Accelerator Energies snapshot
  • 278875 06/04/09 04:03 ejensen Hall A tools screen
  • 278856 06/04/09 01:39 pkpan NMR 64.9%
  • 278841 06/04/09 00:01 S. Sirca Swing shift SUMMARY
  • Wednesday

  • 278839 06/03/09 23:46 S. Sirca New Compton run #17175, HAPPEX run #31470
  • 278834 06/03/09 23:42 S. Sirca Spot++ with run #3166 with He3 target
  • 278829 06/03/09 23:37 miham Target lasers turned OFF/ON
  • 278828 06/03/09 23:36 S. Sirca Spot++ with run #3165 after beam returns
  • 278825 06/03/09 23:13 H Lu Move target to Optics
  • 278824 06/03/09 21:37 S. Sirca Down due to power supply trip (ACC)
  • 278815 06/03/09 20:36 H. Lu NMR 64.9%
  • 278812 06/03/09 20:31 H Lu Spot++ with run #3162
  • 278811 06/03/09 20:19 auto_snap Target Field and Oven Temps snapshot
  • 278810 06/03/09 20:19 auto_snap Target Cell Temps snapshot
  • 278809 06/03/09 20:19 auto_snap Radiation and Ozone snapshot
  • 278808 06/03/09 20:19 auto_snap Happex Charge Asym snapshot
  • 278807 06/03/09 20:19 auto_snap Hall A Tools snapshot
  • 278806 06/03/09 20:19 auto_snap Compton Polarimeter snapshot
  • 278805 06/03/09 20:19 auto_snap BigBox Magnet snapshot
  • 278804 06/03/09 20:18 auto_snap Beam Position and Current snapshot
  • 278803 06/03/09 20:18 auto_snap Accelerator Energies snapshot
  • 278802 06/03/09 20:04 S. Sirca All DAQs recovered
  • 278799 06/03/09 19:49 S. Sirca We're down with computer problems
  • 278797 06/03/09 19:19 S. Sirca Dead time for run #3161 from replay
  • 278796 06/03/09 19:16 R. Michaels spare "hatsv3" available
  • 278790 06/03/09 18:45 miham Target Moved to 3He Cell
  • 278789 06/03/09 18:43 S. Sirca Target to polarized He3
  • 278786 06/03/09 18:11 miham Reference cell filled with N2 (correctly formated)
  • 278785 06/03/09 18:10 miham Reference cell filled with N2
  • 278780 06/03/09 17:49 (DSM) Doug & Pol. 3He Data vs 2 PSI Pure Nitrogen
  • 278779 06/03/09 17:32 miham Spot++ check (#3158) - raster is 4x4mm
  • 278778 06/03/09 17:27 miham Scaler snap shots (#3158) - N2 at 2psig
  • 278773 06/03/09 17:18 Higinbotham Empty Cell 2x2 Raster Plots
  • 278771 06/03/09 17:17 Higinbotham Empty Cell Reactz Now For 2x2 Raster
  • 278765 06/03/09 17:09 V. Sulkosky Updated R-HRS Pedestals for Shower
  • 278762 06/03/09 17:00 miham Spot++ check (#3155) - Raster (2x2)mm
  • 278757 06/03/09 16:49 S. Sirca Reactz with empty reference cell
  • 278756 06/03/09 16:20 miham Reference cell EMPTY (No Gas )
  • 278745 06/03/09 16:08 ellie Day Shift Summary
  • 278739 06/03/09 15:37 ellie Beam has been Restored
  • 278738 06/03/09 15:27 V. Sulkosky Pedestal runs: 3149 and 22021
  • 278737 06/03/09 15:16 ellie Beam Down due to MCC magnet problem
  • 278734 06/03/09 15:01 ellie Reference Cell = D2
  • 278725 06/03/09 14:51 Katich Target moved to Reference Cell Position
  • 278722 06/03/09 14:31 ellie Stated New Compton Run # 17173
  • 278721 06/03/09 14:15 R. Subedi New database directory 20090602 created
  • 278716 06/03/09 13:58 R. Subedi RHRS shower new database after run 22019
  • 278711 06/03/09 13:03 ellie VDC Efficiency Screen Grabs
  • 278706 06/03/09 12:55 ellie RHRS PS6 Changed back to 65535
  • 278705 06/03/09 12:37 ellie Hall A Tools Screen Grab
  • 278704 06/03/09 12:35 ellie Spot++ Screen Grabs
  • 278699 06/03/09 12:18 auto_snap Target Field and Oven Temps snapshot
  • 278698 06/03/09 12:18 auto_snap Target Cell Temps snapshot
  • 278697 06/03/09 12:18 auto_snap Radiation and Ozone snapshot
  • 278696 06/03/09 12:18 auto_snap Happex Charge Asym snapshot
  • 278695 06/03/09 12:18 auto_snap Hall A Tools snapshot
  • 278694 06/03/09 12:18 auto_snap Compton Polarimeter snapshot
  • 278693 06/03/09 12:18 auto_snap BigBox Magnet snapshot
  • 278692 06/03/09 12:18 auto_snap Beam Position and Current snapshot
  • 278691 06/03/09 12:18 auto_snap Accelerator Energies snapshot
  • 278690 06/03/09 12:06 X. Zheng changed PS6 to 10 on right DAQ (clock)
  • 278685 06/03/09 11:41 ellie Beam has been not so good this morning
  • 278684 06/03/09 11:33 ellie Updated LHRS Golden Run to 3142
  • 278683 06/03/09 11:33 ellie Updated LHRS Golden Run to 3142
  • 278682 06/03/09 11:25 ellie RHRS Replay Problems, Run 22015
  • 278681 06/03/09 11:04 ellie Started New Compton Run # 17170
  • 278680 06/03/09 11:03 ellie Started New HAPPEX Run # 31468
  • 278677 06/03/09 11:00 ellie Beam Restored
  • 278676 06/03/09 10:43 gjin BB.theta changed into 75 degrees in db_run.dat since May 12 access
  • 278675 06/03/09 10:38 Katich NMR = 62.53%
  • 278674 06/03/09 09:59 V. Sulkosky Re: Pion Rejector Plots
  • 278673 06/03/09 09:46 ellie Beam Down, Beam Up, Beam Down again
  • 278670 06/03/09 09:43 V. Sulkosky Re: Set Left HRS golden run to run 3128
  • 278657 06/03/09 08:11 ellie Started New Compton Run # 17169
  • 278655 06/03/09 08:05 Boeglin Shift Summary
  • 278640 06/03/09 06:18 pkpan NMR 63%
  • 278637 06/03/09 05:51 ejensen Scrap run #3133
  • 278630 06/03/09 05:15 ejensen Hall A tools screen
  • 278629 06/03/09 05:12 ejensen SPOT++ check
  • 278620 06/03/09 04:18 auto_snap Target Field and Oven Temps snapshot
  • 278619 06/03/09 04:18 auto_snap Target Cell Temps snapshot
  • 278618 06/03/09 04:18 auto_snap Radiation and Ozone snapshot
  • 278617 06/03/09 04:18 auto_snap Happex Charge Asym snapshot
  • 278616 06/03/09 04:18 auto_snap Hall A Tools snapshot
  • 278615 06/03/09 04:17 auto_snap Compton Polarimeter snapshot
  • 278614 06/03/09 04:17 auto_snap BigBox Magnet snapshot
  • 278613 06/03/09 04:17 auto_snap Beam Position and Current snapshot
  • 278612 06/03/09 04:17 auto_snap Accelerator Energies snapshot
  • 278607 06/03/09 03:31 A. Camsonne Set Left HRS golden run to run 3128
  • 278602 06/03/09 03:05 Boeglin Wrong left arm momentum in title for run 3127 and 3128
  • 278595 06/03/09 02:20 pkpan NMR 62.3%
  • 278594 06/03/09 02:19 pkpan NMR 62.3%
  • 278587 06/03/09 00:24 miham Reference cell filled with Deuterium - Deuterium bottle is now EMPTY
  • 278582 06/03/09 00:15 S. Sirca dp vs. phi plot for N2 elastic (run #3122)
  • 278579 06/03/09 00:03 S. Sirca Swing shift SUMMARY
  • Tuesday

  • 278571 06/02/09 23:01 S. Sirca Spot++ on N2 reference cell (run #3122)
  • 278570 06/02/09 22:55 S. Sirca Raw rates with N2 reference cell 21psig at 10uA
  • 278568 06/02/09 22:53 A. Camsonne Compton checks
  • 278566 06/02/09 22:50 S. Sirca Target to N2 reference cell 20psig
  • 278565 06/02/09 22:48 miham NMR Frequency Sweep ( New Polarization is 60.24%) and Absorption spectrum
  • 278564 06/02/09 22:43 A. Camsonne Compton electron detector in
  • 278561 06/02/09 22:12 Shahinyan Target
  • 278556 06/02/09 21:50 S. Sirca New HAPPEX run #31465
  • 278551 06/02/09 21:30 S. Sirca Run #3118 is junk
  • 278550 06/02/09 21:12 S. Sirca Target strip-charts gone
  • 278549 06/02/09 21:09 S. Sirca Left HRS CODA died
  • 278545 06/02/09 20:35 R. Subedi RHRS shower histograms
  • 278544 06/02/09 20:28 Ramesh/Xiaoc Access activity [PVDIS work]
  • 278543 06/02/09 20:17 auto_snap Target Field and Oven Temps snapshot
  • 278542 06/02/09 20:17 auto_snap Target Cell Temps snapshot
  • 278541 06/02/09 20:17 auto_snap Radiation and Ozone snapshot
  • 278540 06/02/09 20:17 auto_snap Happex Charge Asym snapshot
  • 278539 06/02/09 20:17 auto_snap Hall A Tools snapshot
  • 278538 06/02/09 20:17 auto_snap Compton Polarimeter snapshot
  • 278537 06/02/09 20:17 auto_snap BigBox Magnet snapshot
  • 278536 06/02/09 20:17 auto_snap Beam Position and Current snapshot
  • 278535 06/02/09 20:17 auto_snap Accelerator Energies snapshot
  • 278534 06/02/09 20:12 shahinyan NMR 61.79 _jun-2-09
  • 278513 06/02/09 17:53 S. Sirca He3 elastic counts
  • 278508 06/02/09 17:25 S. Sirca New HAPPEX run #31464, BLumi7 saturating
  • 278503 06/02/09 16:42 R. Michaels pre-June-18 HAPPEX work
  • 278491 06/02/09 16:19 ellie Day Shift Summary
  • 278483 06/02/09 16:14 shahinyan Target
  • 278482 06/02/09 16:14 S. Sirca Spot++ with 6x6 raster (10x8 MCC) 2uA run 3107
  • 278476 06/02/09 16:07 L. Zhu target: NMR P=58.6%
  • 278475 06/02/09 16:03 L. Zhu target: to he3
  • 278468 06/02/09 15:31 gjin changed BB angle in db_run.dat
  • 278463 06/02/09 15:06 L. Zhu target: to BeO
  • 278462 06/02/09 14:39 L. Zhu target: he3 to empty
  • 278458 06/02/09 14:29 ellie Changing LHRS Momentum to 2.436 GeV
  • 278457 06/02/09 14:24 A. Camsonne Compton electron detector to garage position for beam tune
  • 278454 06/02/09 12:40 Higinbotham 3He Elastics from Run 3067
  • 278453 06/02/09 12:33 A. Camsonne Added fan for hatsv3
  • 278452 06/02/09 12:17 auto_snap Target Field and Oven Temps snapshot
  • 278451 06/02/09 12:17 auto_snap Target Cell Temps snapshot
  • 278450 06/02/09 12:17 auto_snap Radiation and Ozone snapshot
  • 278449 06/02/09 12:16 auto_snap Happex Charge Asym snapshot
  • 278448 06/02/09 12:16 auto_snap Hall A Tools snapshot
  • 278447 06/02/09 12:16 auto_snap Compton Polarimeter snapshot
  • 278446 06/02/09 12:16 auto_snap BigBox Magnet snapshot
  • 278445 06/02/09 12:16 auto_snap Beam Position and Current snapshot
  • 278444 06/02/09 12:16 auto_snap Accelerator Energies snapshot
  • 278443 06/02/09 12:04 Yi Qiang NMR Cross Calibration
  • 278442 06/02/09 11:36 L. Zhu laser off for tests
  • 278441 06/02/09 10:36 R. Michaels re: sfi2.crl issues (resolved, I think)
  • 278438 06/02/09 10:24 V. Sulkosky re: changes to sfi2.crl
  • 278435 06/02/09 09:47 A. Camsonne Warning message message while fully replaying tun 3069
  • 278432 06/02/09 09:30 L. Zhu target:: laser spectrum
  • 278431 06/02/09 08:21 ellie Hall is now in Controlled Access with sweep completed
  • 278430 06/02/09 08:04 eljensen Owl Shift Summary
  • 278428 06/02/09 07:35 R. Michaels re: changes to sfi2.crl
  • 278427 06/02/09 07:18 ejensen ROC sync check a bit crooked
  • 278418 06/02/09 06:17 pkpan NMR 62.8%
  • 278417 06/02/09 06:17 ejensen SPOT++ check
  • 278414 06/02/09 06:10 ejensen HV screenshot
  • 278413 06/02/09 06:08 ejensen Alarm Handler Screenshot
  • 278412 06/02/09 06:05 ejensen Hall A tools screenshot
  • 278409 06/02/09 05:44 ejensen New Happex Run
  • 278396 06/02/09 04:16 auto_snap Target Field and Oven Temps snapshot
  • 278395 06/02/09 04:16 auto_snap Target Cell Temps snapshot
  • 278394 06/02/09 04:16 auto_snap Radiation and Ozone snapshot
  • 278393 06/02/09 04:16 auto_snap Happex Charge Asym snapshot
  • 278392 06/02/09 04:16 auto_snap Hall A Tools snapshot
  • 278391 06/02/09 04:16 auto_snap Compton Polarimeter snapshot
  • 278390 06/02/09 04:16 auto_snap BigBox Magnet snapshot
  • 278389 06/02/09 04:16 auto_snap Beam Position and Current snapshot
  • 278388 06/02/09 04:16 auto_snap Accelerator Energies snapshot
  • 278387 06/02/09 04:15 ejensen Restarted compton run
  • 278368 06/02/09 02:21 eip changes in the LHRS replay
  • 278367 06/02/09 02:01 ejensen Started new compton run
  • 278364 06/02/09 01:59 pkpan NMR 61.7%
  • 278361 06/02/09 01:37 eip changes in the LHRS replay
  • 278356 06/02/09 01:02 ejensen Restarted Happex
  • 278347 06/02/09 00:06 S. Sirca Swing shift SUMMARY
  • 278341 06/02/09 00:03 flay Target Y & ReactZ (Run# 3084)
  • 278340 06/02/09 00:00 A. Camsonne Odd data for this slot
  • Monday

  • 278339 06/01/09 23:55 S. Sirca Average 11 beam trips/hour
  • 278334 06/01/09 23:10 flay Run# 3083/21956 are junk runs
  • 278321 06/01/09 22:26 S. Sirca roc2 slot16 undefined module in data WARNING during replay
  • 278320 06/01/09 22:18 flay Tool Screen Snapshot (22:20)
  • 278319 06/01/09 22:17 S. Sirca Spot++ for first production run with HRSL at 12.5 deg
  • 278318 06/01/09 22:15 Shahinyan NMR
  • 278317 06/01/09 22:04 S. Sirca Raw rates with polarized He3 cell at 10uA
  • 278312 06/01/09 21:56 S. Sirca Left HRS momentum to 2.320 GeV/c
  • 278309 06/01/09 21:35 shahinyan Target
  • 278306 06/01/09 21:27 S. Sirca Raw rates with N2 ref cell at 20psig and 10uA beam
  • 278305 06/01/09 21:03 gjin target moved to reference N2 cell before the current run
  • 278302 06/01/09 20:56 S. Sirca Start new HAPPEX run #31461
  • 278301 06/01/09 20:55 A. Camsonne Compton EPICS IOC rebooted
  • 278299 06/01/09 20:47 S. Sirca Raw rates with 10 uA on D2 reference cell
  • 278298 06/01/09 20:34 S. Sirca Raster sizes, rates, BPMs with D2 reference cell
  • 278294 06/01/09 20:16 auto_snap Hall A Tools snapshot
  • 278293 06/01/09 20:15 auto_snap Compton Polarimeter snapshot
  • 278292 06/01/09 20:15 auto_snap BigBox Magnet snapshot
  • 278291 06/01/09 20:15 auto_snap Beam Position and Current snapshot
  • 278289 06/01/09 20:15 S. Sirca Spot++ for D2 ref cell run
  • 278290 06/01/09 20:15 auto_snap Accelerator Energies snapshot
  • 278285 06/01/09 19:36 shahinyan NMR
  • 278282 06/01/09 19:20 S. Sirca Raw rates with 11uA
  • 278281 06/01/09 19:16 shahinyan Target
  • 278280 06/01/09 19:13 S. Sirca Run #3067 is He3 elastic with excessive DAQ rates
  • 278275 06/01/09 18:50 S. Sirca CODA unstable, restarted
  • 278246 06/01/09 17:25 shahinyan Target
  • 278245 06/01/09 17:25 S. Sirca Spot++ check
  • 278240 06/01/09 17:03 Jin Huang EPR Calibration for Longitudinal- Field of Cell Moss; Correct Coils has been ON; BB @ 75Degree
  • 278239 06/01/09 16:55 Dezern Movement of spectrometer and B.B.
  • 278238 06/01/09 16:51 S. Sirca BB is at 75 deg and 518A
  • 278237 06/01/09 16:49 S. Sirca Left HRS at 12.5 deg and 2400 MeV/c
  • 278234 06/01/09 16:15 A. Camsonne Power cycled right arm crates
  • 278233 06/01/09 16:07 ellie Day Shift Summary
  • 278231 06/01/09 15:43 A. Camsonne Hall A full survey complete
  • 278230 06/01/09 15:43 ellie Hall A Survey Complete, now in Controlled Access
  • 278228 06/01/09 15:40 Katich NMR 63%
  • 278227 06/01/09 15:02 ellie Called MCC so that we can go into Controlled Access
  • 278226 06/01/09 15:02 adaq EB problems for right HRS
  • 278223 06/01/09 14:58 ellie RHRS now having problems that LHRS had before with ER1
  • 278218 06/01/09 14:48 ellie RHRS DAQ crashed, restarted
  • 278209 06/01/09 14:10 ellie Fixed: adaql1 Computer (Where BB HV GUI is) is not allowing terminals to open
  • 278204 06/01/09 13:41 ellie MCC Called, they are restoring beam
  • 278201 06/01/09 13:11 ellie MCC Called, they're having problems and are taking beam away for a bit
  • 278196 06/01/09 12:54 Boeglin HallA Tools Screen
  • 278195 06/01/09 12:53 Boeglin beam spot size
  • 278194 06/01/09 12:51 ellie Beam has been pretty terrible the past hour
  • 278193 06/01/09 12:18 ellie Started New Compton Run # 17118
  • 278192 06/01/09 12:18 Katich Oven Temperature Snapshot
  • 278191 06/01/09 12:16 ellie It's all good
  • 278190 06/01/09 12:15 auto_snap Target Field and Oven Temps snapshot
  • 278189 06/01/09 12:15 auto_snap Target Cell Temps snapshot
  • 278188 06/01/09 12:15 auto_snap Radiation and Ozone snapshot
  • 278187 06/01/09 12:15 auto_snap Happex Charge Asym snapshot
  • 278186 06/01/09 12:15 auto_snap Hall A Tools snapshot
  • 278185 06/01/09 12:15 auto_snap Compton Polarimeter snapshot
  • 278184 06/01/09 12:15 auto_snap BigBox Magnet snapshot
  • 278183 06/01/09 12:15 auto_snap Beam Position and Current snapshot
  • 278182 06/01/09 12:15 auto_snap Accelerator Energies snapshot
  • 278181 06/01/09 12:11 Katich Target Cell Temperature Snapshot
  • 278180 06/01/09 12:10 ellie Everything on the DAQ computer rebooted, started, and up and running
  • 278173 06/01/09 12:05 ellie Getting everything started from Cold Start
  • 278172 06/01/09 11:55 ellie Attempted to Reboot ER1. Accidentally rebooted entire system on the DAQ computers
  • 278170 06/01/09 11:52 ellie LHRS Still not working, ER1 Problem Details
  • 278165 06/01/09 11:33 adaq problems with ER1 for left HRS and dead time
  • 278164 06/01/09 11:32 ellie LHRS DAQ Not fixed from Reboot
  • 278163 06/01/09 11:30 ellie adaql1 Computer (Where BB HV GUI is) is not allowing terminals to open
  • 278160 06/01/09 11:24 ellie LHRS Deadtime high, ROCs giving error messages, rebooting it all
  • 278156 06/01/09 11:19 Katich NMR 60.9%
  • 278150 06/01/09 10:27 ellie Started New HAPPEX Run # 31460
  • 278145 06/01/09 10:13 ellie RHRS Q1 Red Helium Alerts during Day Shift, Should now be over
  • 278143 06/01/09 10:13 ellie RHRS Q1 Red Helium Alerts during Day Shift, Should now be over
  • 278137 06/01/09 09:43 ellie Start New Compton Run # 17117
  • 278122 06/01/09 08:00 miham OWL Shift Summary (06/01/09)
  • 278112 06/01/09 06:50 pkpan NMR_63%
  • 278096 06/01/09 04:52 miham Left arm coda crash
  • 278084 06/01/09 04:15 auto_snap Target Field and Oven Temps snapshot
  • 278083 06/01/09 04:14 auto_snap Target Cell Temps snapshot
  • 278082 06/01/09 04:14 auto_snap Radiation and Ozone snapshot
  • 278081 06/01/09 04:14 auto_snap Happex Charge Asym snapshot
  • 278080 06/01/09 04:14 auto_snap Hall A Tools snapshot
  • 278079 06/01/09 04:14 auto_snap Compton Polarimeter snapshot
  • 278078 06/01/09 04:14 auto_snap BigBox Magnet snapshot
  • 278077 06/01/09 04:14 auto_snap Beam Position and Current snapshot
  • 278076 06/01/09 04:14 auto_snap Accelerator Energies snapshot
  • 278075 06/01/09 04:07 miham Again Double Alarm
  • 278066 06/01/09 03:14 miham New Compton run #17114 and its problems
  • 278065 06/01/09 02:57 miham Double alarm
  • 278064 06/01/09 02:54 miham HAPPEX is working again
  • 278057 06/01/09 02:16 pkpan NMR_62%
  • 278054 06/01/09 01:51 miham HAPPEX and LUMI not working
  • 278049 06/01/09 01:12 miham New Compton Run #17112 - Compton not working
  • 278048 06/01/09 01:02 miham Spot++ Check (#3031)
  • 278043 06/01/09 00:32 miham Half Wave plate Inserted (now IN)
  • 278034 06/01/09 00:17 miham Hall-A tools and Alarms
  • 278033 06/01/09 00:16 miham HV GUI Screen shot - All green
  • 278032 06/01/09 00:00 adaq Shift Summary